### v1

# Seeking Consistent Flat Minima for Better Domain Generalization via Refining Loss Landscapes

Abstract
Domain generalization aims to learn a model from multiple training domains and generalize it to unseen test domains. Recent theory has shown that seeking the deep models, whose parameters lie in the flat minima of the loss landscape, can significantly reduce the out-of-domain generalization error. However, existing methods often neglect the consistency of loss landscapes in different domains, resulting in models that are not simultaneously in the optimal flat minima in all domains, which limits their generalization ability. To address this issue, this paper proposes an iterative Self-Feedback Training (SFT) framework to seek consistent flat minima that are shared across different domains by progressively refining loss landscapes during training. It alternatively generates a feedback signal by measuring the inconsistency of loss landscapes in different domains and refines these loss landscapes for greater consistency using this feedback signal. Benefiting from the consistency of the flat minima within these refined loss landscapes, our SFT helps achieve better out-of-domain generalization. Extensive experiments on DomainBed demonstrate superior performances of SFT when compared to state-of-the-art sharpness-aware methods and other prevalent DG baselines. On average across five DG benchmarks, SFT surpasses the sharpness-aware minimization by 2.6% with ResNet-50 and 1.5% with ViT-B/16, respectively. The code will be available soon.
1 Introduction
The task of Domain Generalization (DG) is to learn a model from multiple training domains so that it can generalize well to unseen test domains [8]. Though modern deep learning has achieved remarkable success in many areas [3, 4, 5, 6], it assumes that training and test data are independent and identically distributed (IID). This assumption is often violated in real applications, which significantly degrades the performance of deep models [8]. To address this problem, abundant domain generalization (DG) algorithms have been developed from different perspectives, e.g., invariant or causal representation learning [12, 50, 13], disentangled representation learning [53, 54], distributionally robust optimization [14], meta-learning [15, 16, 17], data augmentation [20, 21], etc. Nonetheless, it is still an open problem for the communities of modern machine learning.
((a))
((b))
Figure 1: Loss Landscapes without and with consistency. The left subplot (a) illustrates the inconsistency of loss landscapes across different domains, which arise due to domain shifts. This paper proposes refining these landscapes to achieve improved consistency, as demonstrated in the right subplot (b).
Recently, the loss landscape [24], which refers to the shape of a loss function within the parameter space, provides a unified perspective for understanding various DG algorithms. Essentially, the training process of deep models is to seek the minimum point within the training loss landscape. From the view of loss landscapes [24], the ultimate goal of various DG methods is to construct a suitable loss landscape so that the loss minimum point sought on training domains is exactly the loss minimum point on unseen test domains. In fact, significant studies have revealed the connection between the geometry of the loss landscape (particularly the flatness of minima [25]) and generalization from both theoretical and empirical perspectives [27, 28, 29, 34]. This connection has shown potential in enabling novel methods to model training that yield better generalization [34]. For example, by penalizing the sharpness of the loss landscape, numerous methods [34, 31, 32] have exhibited superior generalization on the DomainBed benchmark [11] and many other datasets [35, 36]. However, in the DG scenario, the domain shifts contribute to the drastic discrepancy of loss landscapes [70], and thus loss landscapes in test domains may not exhibit similar geometries as in training domains. Consequently, the flat minima sought on the training domains do not have a low loss value in the unseen test domains, meaning that the learned models may have inferior performance in the test domains. Therefore, maintaining consistency of the loss landscapes between the training and unseen test domains is crucial for better generalization, as it enables us to seek the consistent flat minima that perform well across all domains.
To address this issue, this paper proposes refining the loss landscapes to maintain their consistency in different domains. However, there are several primary challenges to be addressed. Firstly, in the DG setting, the inaccessibility of the test domain restricts us to using limited training domains for refining loss landscapes. Then, how to ensure that the loss landscape consistency achieved on training domains can transfer to unseen test domains becomes a challenging endeavor. Secondly, refining the loss landscapes usually means high complexity. Typically, it involves the selection of different model architectures [75, 73, 74, 76] or the artificial design of novel loss functions [50, 47, 46]. Architecture selection suffers from substantial computational overhead [77], while artificially designing an effective loss function for DG is also exceedingly intractable. Therefore, finding efficient approaches to refine various training loss landscapes is far from straightforward.
Inspired by studies on the self-refinement of large-scale language models [1, 2], this paper introduces an iterative two-phase framework called Self-Feedback Training (SFT) to progressively refine loss landscapes and find consistent flat minima across different domains. Specifically, it alternatively generates a feedback signal by measuring the inconsistency of loss landscapes in different domains in the feedback phase, and then refines these loss landscapes for greater consistency using this feedback signal in the refinement phase. To ensure the consistency of loss landscapes between the training and unseen test domains, this paper introduces a training-domain split scheme in the feedback phase. In each iteration, the SFT quantifies the inconsistency by comparing the loss landscapes of the trained and held-out domains. Theoretical analysis (in the supplementary file) shows that in this way, the sharpness of the test loss can be constrained by that of the training domains, which also means the transferability of the loss landscapes consistency. Furthermore, to enable low-complexity refinement of loss landscapes, a landscape refiner is introduced in the refinement phase. Instead of altering the model architecture or devising novel loss functions, the refiner continuously modifies the geometries of the loss landscapes by generating dynamical soft labels as training progresses. Compared to existing methods [75, 76], our landscape refining method not only substantially reduces the computational overhead, but also overcomes the over-confident issue brought by traditional one-hot labels [39]. To our best knowledge, this is the first work to refine the loss landscapes from a label-based perspective. Benefiting from the consistency of the flat minima within these refined loss landscapes, our SFT helps achieve better out-of-domain generalization. Extensive experiments on a synthetic dataset and five popular DG benchmarks demonstrate that SFT achieves superior performance compared to popular sharpness-aware methods [34, 103, 96, 104, 95] and other prevalent DG methods.
In summary, our key contributions are as follows:
  • A novel self-feedback training framework, which could refine loss landscapes and find consistent flat minima across different domains, is proposed to obtain models with better domain generalization.
  • A landscape refiner is introduced to facilitate the efficient refinement of loss landscapes. To our best knowledge, it is the first work to refine the loss geometry from the perspective of labels.
  • Extensive experiments on the synthetic dataset and five popular DG benchmarks demonstrate the superiority of SFT over popular sharpness-aware methods and other prevalent DG methods.
2 Related Work
In this section, we review related areas of research: studies on loss landscapes for model generalization and domain generalization.
Loss landscapes for model generalization. The idea of searching for “flat” minima of loss landscapes can be traced back to Hochreiter and Schmidhuber [26], who proposed penalizing the sharpness of minima as a form of regularization, related to the Minimum Description Length (MDL) principle. Over the past few decades, numerous studies [27, 28, 29] have linked the flatness of loss landscapes to the generalization ability of deep models. As a result, many optimization methods have been developed to find flatter minima, such as Entropy-SGD [32] and the diffusion approach [31]. More recently, two methods, Stochastic Weight Averaging (SWA) [71] and Sharpness-Aware Minimization (SAM) [34], have gained significant attention for their effectiveness and scalability. SWA finds flatter minima by averaging model parameters across the optimization trajectory, while SAM seeks parameters within regions of the loss landscape that exhibit uniformly low loss values. Jean Kaddour et al.  [72] provided a detailed comparison of these methods, showing that SAM tends to lead models to wider basins, while SWA merely helps locate the center of one certain basin. In this paper, we focus on addressing the inconsistency of loss landscapes when SAM is applied in domain generalization (DG) scenarios.
Loss landscapes for domain generalization. Recently, optimization techniques aiming at finding flat minima in loss landscapes have also been applied to improve the out-of-domain generalization of deep models [30]. Junbum Cha et al. [70] introduced the flatness to the domain generalization and proposed a novel Stochastic Weight Averaging Densely (SWAD), which is an extension of SWA with a dense and overfit-aware stochastic weight sampling strategy. Zhang et al. [103] pointed out that SAM-based methods are limited to zeroth-order flatness, which may be insufficient to discriminate minima with low generalization error. Thus, they developed a novel method named GAM, which targets first-order flatness to enhance generalization. Moreover, FAD [104] optimizes both zeroth- and first-order flatness simultaneously for domain generalization. However, these methods primarily focus on seeking better flat minima for DG, ignoring the influence of landscape discrepancy across multiple domains. In contrast, Giambattista Parascandolo et al. [56] considered the domain discrepancy via formalizing a notion of consistency for minima of loss surfaces, and then proposed an algorithm (ANDMask) to alter the optimization trajectory through masking inconsistent gradient components among training domains. Different from all of the above methods that consider loss landscapes to be static, we attempt to refine the loss surfaces dynamically for consistency and finally seek a consistent flat minima across various domains.
3 Methodology
3.1 Problem Formulation and Preliminaries
Let us begin with a formal description of domain and domain generalization (DG). Let $\mathcal{X}$ and $\mathcal{Y}$ denote the input sample space and the category space, respectively. In the DG problem, we use data sampled from $p$ training distributions $\{\mathcal{D}_{d}\}_{d=1}^{p}$ and $q$ test distributions $\{\mathcal{D}_{d}\}_{d=p+1}^{p+q}$, each of which defines on the joint space $\mathcal{X}\times\mathcal{Y}$. $D_{d}=\{(\boldsymbol{x}_{i}^{(d)},y_{i}^{(d)})\}_{i=1}^{n_{d}}$ denotes the dataset sampled from the $d$-th distribution $\mathcal{D}_{d}$, which is referred to as the $d$-th domain. $(\boldsymbol{x}_{i}^{(d)},y_{i}^{(d)})\in\mathcal{X}\times\mathcal{Y}$ denotes the $i$-th sample from the domain $\mathcal{D}_{d}$ and $n_{d}$ denotes the number of samples in the $d$-th domain. Notably, each domain contains different domain statistics but shares the same category space. Domain generalization aims to train a model $\boldsymbol{f}_{\boldsymbol{\theta}}:\mathcal{X}\rightarrow\mathcal{Y}$ on the $p$ training domains so that it can generalize well to the $q$ novel target domains. Please note that the $q$ test domains are inaccessible during training, which differs from the problem of domain adaptation [83].
In this paper, we consider the model mentioned above to be a parametric deep neural network $\boldsymbol{f}_{\boldsymbol{\theta}}$, with $\boldsymbol{\theta}$ denoting the parameters of the network. Standard ERM training of networks usually utilizes the cross entropy (CE) as the loss function:
$\mathcal{L}_{D_{tr}}^{\mathrm{CE}}(\boldsymbol{\theta})=-\sum_{d=1}^{p}\sum_{i=1}^{n_{d}}\mathbf{y}_{i}^{(d)\mathbf{T}}\log\boldsymbol{f}_{\boldsymbol{\theta}}(\boldsymbol{x}_{i}^{(d)}),$ (1)
where $\mathcal{L}_{D_{tr}}^{\mathrm{CE}}$ represents the cross-entropy loss computed on the training set, which is an aggregation of samples from all the training domains. In this paper, we use the superscript on $\mathcal{L}$ to indicate the type of loss function (e.g., cross entropy), while the subscript denotes the dataset on which the loss is calculated. $\boldsymbol{f}_{\boldsymbol{\theta}}(\boldsymbol{x}_{i}^{(d)})$ is a vector with $N$ dimensions (where $N$ is the number of classes), and the $j$-th entry of the vector represents the predictive probability that $\boldsymbol{x}_{i}^{(d)}$ belongs to the $j$-th class. $\mathbf{y}_{i}^{(d)}$ denotes the one-hot encoded label of $y_{i}^{(d)}$, whose superscript “$\mathbf{T}$” indicates the matrix transpose operation.
However, optimizing the training loss value only (e.g., ERM) can easily lead to suboptimal model quality [34]. Thus, SAM [34] tries to simultaneously minimize loss value and loss sharpness. It defines the sharpness by measuring how quickly the training loss can be increased by adding a small perturbation $\boldsymbol{\epsilon}$ (with norm less than $\rho$) to the current parameter $\boldsymbol{\theta}$. To avoid sharpness being dependent on the perturbation $\boldsymbol{\epsilon}$, the worst case is considered when evaluating sharpness:
$\mathcal{L}_{D_{tr}}^{\mathrm{CES}}(\boldsymbol{\theta})=\max_{||\boldsymbol{\epsilon}||\leq\rho}\mathcal{L}_{D_{tr}}^{\mathrm{CE}}(\boldsymbol{\theta}+\boldsymbol{\epsilon})-\mathcal{L}_{D_{tr}}^{\mathrm{CE}}(\boldsymbol{\theta}).$ (2)
Here, we use $\mathcal{L}_{D_{tr}}^{\mathrm{CES}}$ to denote the sharpness measure for $\mathcal{L}_{D_{tr}}^{\mathrm{CE}}$. Since the perturbation strength $\rho$ is small enough, the above optimization could be solved approximately via Taylor expansion, leading to the optimal perturbation $\bar{\boldsymbol{\epsilon}}_{D_{tr}}^{\mathrm{CE}}$ for $\mathcal{L}_{D_{tr}}^{\mathrm{CE}}$ as:
$\bar{\boldsymbol{\epsilon}}_{D_{tr}}^{\mathrm{CE}}(\boldsymbol{\theta})=\rho\frac{\nabla_{\boldsymbol{\theta}}\mathcal{L}_{D_{tr}}^{\mathrm{CE}}(\boldsymbol{\theta})}{||\nabla_{\boldsymbol{\theta}}\mathcal{L}_{D_{tr}}^{\mathrm{CE}}(\boldsymbol{\theta})||}.$ (3)
Then, the final loss function used in SAM combines both the loss value and the loss sharpness, and is given by:
$\mathcal{L}_{D_{tr}}^{\mathrm{SAM}}(\boldsymbol{\theta})=\mathcal{L}_{D_{tr}}^{\mathrm{CE}}(\boldsymbol{\theta}+\bar{\boldsymbol{\epsilon}}_{D_{tr}}^{\mathrm{CE}}).$ (4)
During training with (stochastic) gradient descent, the contribution of $\nabla_{\boldsymbol{\theta}}\bar{\boldsymbol{\epsilon}}_{D_{tr}}^{\mathrm{CE}}(\boldsymbol{\theta})$ could be neglected due to small perturbation strength $\rho$.
1
Input: Training domains $D_{tr}$, learning rate $\eta$ and other hyperparameters.
Output: Model parameter $\boldsymbol{\theta}$.
2
3Init: Parameters of the model $\boldsymbol{f}_{\boldsymbol{\theta}}$ and refiner $\boldsymbol{g}_{\boldsymbol{\phi}}$;
4while not converge do
5
6      The Feedback Phase:
7      Randomly select two training domains $D_{d}$ and $D_{d^{\prime}}$.
8      Update the model parameter $\boldsymbol{\theta}$ on the domain $D_{d}$:
9      $\boldsymbol{\theta}\leftarrow\boldsymbol{\theta}-\eta\nabla_{\boldsymbol{\theta}}\mathcal{L}_{D_{d}}^{\mathrm{SAM-SL}}(\boldsymbol{\theta},\boldsymbol{\phi})$.
10      Calculate the feedback signal $|\mathcal{L}_{D_{d}}^{\mathrm{CES-SL}}(\boldsymbol{\theta},\boldsymbol{\phi})-\mathcal{L}_{D_{d^{\prime}}}^{\mathrm{CES-SL}}(\boldsymbol{\theta},\boldsymbol{\phi})|$ using both domains.
11      The Refinement Phase:
12      Calculate the total loss for refinement, $\mathcal{L}_{D_{tr}}^{\mathrm{Refine}}(\boldsymbol{\theta},\boldsymbol{\phi})$, by leveraging the above feedback signal.
13      Update the refiner parameter: $\boldsymbol{\phi}\leftarrow\boldsymbol{\phi}-\eta\nabla_{\boldsymbol{\phi}}\mathcal{L}_{D_{tr}}^{\mathrm{Refine}}(\boldsymbol{\theta},\boldsymbol{\phi})$.
14 end while
Algorithm 1 SFT for Domain Generalization.
3.2 Self-Feedback Training
To address the issue of landscape inconsistency when applying SAM-based methods to DG scenarios, we introduce an iterative two-phase framework called Self-Feedback Training (SFT) to progressively refine loss landscapes and find consistent flat minima across different domains. Specifically, it alternatively generates a feedback signal by measuring the inconsistency of loss landscapes in different domains in the feedback phase, and then refines these loss landscapes for greater consistency using this feedback signal in the refinement phase. In the following, this paper provides a detailed description of this framework.
3.2.1 The Feedback Phase
In this phase, SFT mainly detects inconsistency between various loss landscapes and generates a feedback signal.
Let’s start by analyzing the key elements necessary for obtaining a feedback signal. First, due to the absence of test domains, we are limited to assessing the landscape inconsistency within the training domains. Making full use of the training domains to simulate domain shifts that may occur in unseen domains will help improve the model’s generalization ability to those unknown domains. Second, the training loss should depend on some learnable parameters, which enables us to optimize them for landscape refinement in the next phase. Finally, we need to determine which measurement to use when assessing the landscape inconsistency.
Considering the above analysis, we choose to implement a training-domain split scheme, where at each iteration, two training domains, $D_{d}$ and $D_{d^{\prime}}$, are randomly selected. $D_{d}$ is used to train the model, while $D_{d^{\prime}}$ is held out to evaluate landscape inconsistency in comparison to $D_{d}$. During training on the domain $D_{d}$, we employ the training loss that depends on the parameters of a landscape refiner $\boldsymbol{g}_{\boldsymbol{\phi}}$. This refiner enables the efficient refinement of loss landscapes via generating soft labels for each sample:
$\tilde{\mathbf{y}}_{i}^{(d)}=\boldsymbol{g}_{\boldsymbol{\phi}}(\boldsymbol{x}_{i}^{(d)}).$ (5)
Then, the loss function for sharpness-aware minimization can be expressed as:
$\mathcal{L}_{D_{d}}^{\mathrm{SAM-SL}}(\boldsymbol{\theta},\boldsymbol{\phi})=\mathcal{L}_{D_{d}}^{\mathrm{CE-SL}}(\boldsymbol{\theta}+\bar{\boldsymbol{\epsilon}}_{D_{d}}^{\mathrm{CE-SL}},\boldsymbol{\phi}).$ (6)
Here, “SL” represents the utilization of soft labels. $\mathcal{L}_{D_{d}}^{\mathrm{CE-SL}}$ and $\bar{\boldsymbol{\epsilon}}_{D_{d}}^{\mathrm{CE-SL}}$ are defined as:
$\mathcal{L}_{D_{d}}^{\mathrm{CE-SL}}(\boldsymbol{\theta},\boldsymbol{\phi})=-\sum_{i=1}^{n_{d}}\tilde{\mathbf{y}}_{i}^{(d)\mathbf{T}}\log\boldsymbol{f}_{\boldsymbol{\theta}}(\boldsymbol{x}_{i}^{(d)}),$ (7)
$\bar{\boldsymbol{\epsilon}}_{D_{d}}^{\mathrm{CE-SL}}(\boldsymbol{\theta},\boldsymbol{\phi})=\rho\frac{\nabla_{\boldsymbol{\theta}}\mathcal{L}_{D_{d}}^{\mathrm{CE-SL}}(\boldsymbol{\theta},\boldsymbol{\phi})}{||\nabla_{\boldsymbol{\theta}}\mathcal{L}_{D_{d}}^{\mathrm{CE-SL}}(\boldsymbol{\theta},\boldsymbol{\phi})||}.$ (8)
For evaluating landscape inconsistency, one direct approach is to look at the difference in loss values across domains. However, the loss value only captures zero-order information. Considering that loss sharpness is closely related to the Hessian matrix and reflects second-order loss information, we instead use the difference in loss sharpness, $|\mathcal{L}_{D_{d}}^{\mathrm{CES-SL}}(\boldsymbol{\theta},\boldsymbol{\phi})-\mathcal{L}_{D_{d^{\prime}}}^{\mathrm{CES-SL}}(\boldsymbol{\theta},\boldsymbol{\phi})|$, to quantify the landscape consistency, which serves as a feedback signal to the landscape refiner. Here, $\mathcal{L}_{D_{d}}^{\mathrm{CES-SL}}$ is define as:
$\mathcal{L}_{D_{d}}^{\mathrm{CES-SL}}(\boldsymbol{\theta},\boldsymbol{\phi})=\mathcal{L}_{D_{d}}^{\mathrm{CE-SL}}(\boldsymbol{\theta}+\bar{\boldsymbol{\epsilon}}_{D_{d}}^{\mathrm{CE-SL}},\boldsymbol{\phi})-\mathcal{L}_{D_{d}}^{\mathrm{CE-SL}}(\boldsymbol{\theta},\boldsymbol{\phi}).$ (9)
Similarly, $\mathcal{L}_{D_{d^{\prime}}}^{\mathrm{CES-SL}}$ can be defined. Please note that the landscape refiner parameter $\boldsymbol{\phi}$ will be optimized in the refinement phase using the feedback signal and could be treated as constant in this phase.
((a))
((b))
((c))
((d))
((e))
((f))
((g))
((h))
Figure 2: 2D visualization of loss surfaces at each domain with/without landscape refinement. The first row shows the inconsistency of loss surfaces using one-hot labels (without refinement); the second row shows the improved landscape consistency using soft labels generated by the landscape refiner. The final well-trained model is marked by “$+$”.
3.2.2 The Refinement Phase
After obtaining the feedback signal, this phase will use it to optimize the landscape refiner to enhance the consistency of different landscapes.
As mentioned earlier, the landscape refiner refines the loss landscapes by generating soft labels for subsequent training. There are several important considerations about these labels. First, it is essential to maintain the correctness of the labels as much as possible, as this directly impacts the effectiveness of training. Second, the soft labels help improve the consistency of the refined loss landscapes, which is central to addressing the landscape inconsistency problem discussed in this paper. Moreover, there are two potential directions to refining the loss landscapes for consistency: making them both sharper or making them both flatter. Given the relationship between flatness and generalization, the soft labels need to be designed to encourage the loss landscapes to become flatter.
To ensure label correctness, we could use the standard cross-entropy loss to force the generated soft labels to approximate one-hot labels. However, this approach carries the risk of overfitting to the one-hot labels, leading to overconfidence and limiting the model’s generalization ability. One technique that attempts to mitigate this issue is label smoothing [39], but it merely introduces noise to the one-hot labels, making them less informative and unable to capture the underlying relationships between categories and different input samples. This paper focuses on a more fundamental problem: minimizing the “distance” between the refiner’s output and the true label space for each category. For instance, if the first category is the true label for the sample $\boldsymbol{x}_{i}$, the corresponding label space is defined as:
$C_{1}=\{(q_{1},\ldots,q_{N})\mid\forall k\neq 1:q_{1}\geq\alpha q_{k},\sum_{k=1}^{N}q_{k}=1\}.$ (10)
where $\alpha\geq 1$ is a hyperparameter that represents the minimum ratio between $q_{1}$ and $q_{k}$. Similar to the concept of “distance” in Euclidean space, we define the “distance” between the output $\tilde{\mathbf{y}}=\boldsymbol{g}_{\boldsymbol{\phi}}(\boldsymbol{x})$ and the label space $C_{1}$ as:
$\min_{\mathbf{y}}\,\mathrm{KL}(\mathbf{y}\,||\,\tilde{\mathbf{y}})\quad\text{subject to}\quad\mathbf{y}\in C_{1}.$ (11)
We propose an efficient algorithm (Algorithm 2) to find the optimal solution $\mathbf{y}^{\star}$ for this problem, which converges in $N$ steps ($N$ is the number of classes) and is computationally more efficient than using common convex programming tools. For the condition that $y_{i}\neq 1$, we can also get the optimal solution via category swapping before and after applying Algorithm 2. Once having $\mathbf{y}^{\star}$, the task becomes minimizing the KL divergence between $\mathbf{y}^{\star}$ and $\tilde{\mathbf{y}}$: $\mathrm{KL}(\mathbf{y}^{\star}\,||\,\tilde{\mathbf{y}})$. Since $\mathbf{y}^{\star}$ depends on $\boldsymbol{\phi}$ and Algorithm 2 may involve non-differentiable operations, we adopt a widely-used alternative iterative strategy: we treat $\mathbf{y}^{\star}$ as constant during each update step for $\boldsymbol{\phi}$, and after updating $\boldsymbol{\phi}$, we recalculate $\mathbf{y}^{\star}$ using Algorithm 2. Minimizing the KL divergence then leads to an objective similar to cross entropy:
$\mathcal{L}_{D_{tr}}^{\mathrm{PCE}}(\boldsymbol{\phi})=-\sum_{d=1}^{p}\sum_{i=1}^{n_{d}}\mathbf{y}_{i}^{\star(d)\mathbf{T}}\log\boldsymbol{g}_{\boldsymbol{\phi}}(\boldsymbol{x}_{i}^{(d)}),$ (12)
which we refer to as projection cross-entropy (PCE) loss. Interestingly, we find that label smoothing is just a special case of our method, applicable only to a small subset of samples. More detailed derivations and further discussion are available in the supplementary file.
Input: The hyperparameter $\alpha$, $\tilde{\mathbf{y}}=(p_{1},\dots,p_{N})$.
Output: The optimal solution $\mathbf{y}^{\star}=(q_{1},\dots,q_{N})$.
1
2Initialization: $A\leftarrow\emptyset$, $B\leftarrow\{j|\alpha p_{j}>p_{1}\}$, $t\leftarrow 1$.
3
4Sort the elements of $B$ in descending order: $p_{j_{1}}\geq\dots\geq p_{j_{|B|}}$.
5
6Update: $A\leftarrow A\cup\{j_{1}\}$, $t\leftarrow t+1$.
7
8while $t\leq|B|$ do
9       if $\left(p_{1}^{\alpha}\left(\prod_{j\in A}\alpha p_{j}\right)\right)^{\frac{1}{|A|+\alpha}}<\alpha p_{j_{t}}$ then
10             Update: $A\leftarrow A\cup\{j_{t}\}$, $t\leftarrow t+1$.
11
12      else
13             Break
14       end if
15
16 end while
17
18Calculate $q_{1}$: $q_{1}=\left(p_{1}^{\alpha}\left(\prod_{j\in A}\alpha p_{j}\right)\right)^{\frac{1}{|A|+\alpha}}$.
19
20for $i\in\{2,\dots,N\}$ do
21       if $i\in A$ then
22             $q_{i}=q_{1}/\alpha$.
23      else
24             $q_{i}=p_{i}$.
25       end if
26
27 end for
28
29Normalize $q_{1},\dots,q_{N}$ such that $\sum_{i}q_{i}=1$.
Algorithm 2 KL Divergence Minimization
To encourage the loss landscapes to become flatter while pursuing consistency, we add the the sharpness term to the previously obtained feedback signal (the sharpness difference) as an extra penalty. Then, the final loss function is defined as:
$\mathcal{L}_{D_{tr}}^{\mathrm{Refine}}(\boldsymbol{\theta},\boldsymbol{\phi})=\mathcal{L}^{\mathrm{PCE}}_{D_{tr}}(\boldsymbol{\phi})+\lambda_{1}\mathcal{L}_{D_{d}}^{\mathrm{CES-SL}}(\boldsymbol{\theta},\boldsymbol{\phi})$
$+\lambda_{2}|\mathcal{L}_{D_{d}}^{\mathrm{CES-SL}}(\boldsymbol{\theta},\boldsymbol{\phi})-\mathcal{L}_{D_{d^{\prime}}}^{\mathrm{CES-SL}}(\boldsymbol{\theta},\boldsymbol{\phi})|.$ (13)
Here, the hyperparameters $\lambda_{1}$ and $\lambda_{2}$ control the relative importance of each term in the loss function. By using this loss function to update the landscape refiner $\boldsymbol{g}_{\boldsymbol{\phi}}$, the generated soft labels could help adjust the loss landscapes, making them more consistent and flatter over time.
As the iterative process of feedback and refinement continues, the model benefits from these refined landscapes, guiding it toward more stable and consistent optimization paths. Ultimately, the model can seek the consistent flat minima across domains, which helps achieve better out-of-domain generalization.
4 Experiments
In this section, we begin by designing a toy experiment to visually demonstrate the landscape consistency achieved through self-feedback training, and explore how the landscape consistency impacts DG performance. A thorough theoretical analysis is provided in the supplementary file. After that, we perform extensive experiments on real-world datasets to validate the effectiveness of the SFT method.
4.1 Experiments on Toy Dataset
Data generation. In our toy experiments, we use a hierarchical Gaussian model. Specifically, we assume that the data from the $i$-th class in the $j$-th domain follows a two-dimensional Gaussian distribution, i.e., $p(\boldsymbol{x})=\mathcal{N}(\boldsymbol{x};\boldsymbol{\mu}_{ij},\boldsymbol{\Sigma}_{ij}),$ where $\boldsymbol{\mu}_{ij}\in\mathbb{R}^{2}$ and $\boldsymbol{\Sigma}_{ij}\in\mathbb{R}^{2\times 2}$ are the mean vector and covariance matrix for the $j$-th domain in the $i$-th class, respectively. The mean vector $\boldsymbol{\mu}_{ij}$ is sampled from another Gaussian distribution: $p(\boldsymbol{\mu}_{ij})=\mathcal{N}(\boldsymbol{\mu}_{ij};\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i}),$ where $\boldsymbol{\mu}_{i}$ and $\boldsymbol{\Sigma}_{i}$ represent the mean and covariance for the $i$-th class across domains. In our setup, we generate a dataset with three classes and four domains. The first three domains are used for training, while the fourth domain serves as the test set. Further details can be found in the supplementary file. A linear classifier is employed in these toy experiments.
Loss landscape visualization. To investigate the consistency of the landscape across the four domains after self-feedback training, we visualize the 2D loss surfaces in Fig.2. We first choose three model weights $(\theta_{1},\theta_{2},\theta_{3})$,111 Here, $\theta_{1}$ represents the parameters of the well-trained model, while $\theta_{2}$ and $\theta_{3}$ correspond to two randomly initialized models. and use them to derive two axes, $e_{1}$ and $e_{2}$, through the Gram–Schmidt process. Then, we compute loss values by varying the coefficients $\beta_{1}$ and $\beta_{2}$ in the linear combination $\theta_{1}+\beta_{1}e_{1}+\beta_{2}e_{2}$, where $\beta_{1},\beta_{2}\in[-2,2]$ in our experiments. From Fig.2, we observe that the loss surfaces using soft labels generated by the landscape refiner exhibit greater consistency across domains compared to those obtained using one-hot labels. Besides, we find that the final model sought by SFT, marked by a “+”, indeed locates the flat minima within the more consistent loss surfaces. These results demonstrate the effectiveness of SFT.
Influence of the landscape consistency on DG performances. In this experiment, we gradually increase the values of $\lambda_{2}$ in (3.2.2) to strengthen the promotion of landscape consistency, while tracking the changes in DG performance. When there’s no ambiguity, we drop the subscript on $\lambda_{2}$ and refer to it simply as $\lambda$. As shown in Fig.3, if the perturbation strength $\rho$ is enough small, the out-of-domain accuracy initially increases and then decreases as $\lambda$ increases (see Fig.3(b)). This phenomenon can be explained as a trade-off between maintaining label correctness and enhancing the landscape consistency: Initially, with a small value of $\lambda$, the soft labels remain accurate, and increasing $\lambda$ helps improve out-of-domain generalization by encouraging landscape consistency. However, when $\lambda$ becomes too large, it compromises the correctness of the soft labels, leading to a sharp drop in accuracy for both the training (see Fig.3(a)) and test domains. Besides, as the perturbation strength $\rho$ increases, the accuracy peak on the test domain shifts to the left, which is because a larger sharpness value will be obtained with larger $\rho$, alleviating the need for excessively large $\lambda$ to promote consistency. Finally, when $\rho$ reaches 0.5, accuracies on the training and test domains deteriorate rapidly, similar to the behavior of the SAM [34].
((a))
((b))
Figure 3: Performances of classification with the varied hyperparameter $\lambda$. The left (a) and right subplots (b) show the performance on the training and test domains, respectively.
4.2 Experiments on Real Dataset
Table 1: Comparison with popular DG methods with ResNet-50 pre-trained on ImageNet. The table reports the average out-of-domain accuracy on five DG datasets. Each out-of-domain performance is an average of three different runs with distinct train-validation splits. We highlight the best results in bold and underline the second best results. Results marked by $\dagger$ and $\ddagger$ are cited from Gulrajani and Lopez-Paz [11] and Cha et al. [70], respectively.
      Algorithms       VLCS       PACS       OfficeHome       TerraIncognita       DomainNet       Avg.
      ERM [45]       77.5$\pm$0.4       85.5$\pm$0.2       66.5$\pm$0.3       46.1$\pm$1.8       40.9$\pm$0.1       63.3
      IRM [50]       78.5$\pm$0.5       83.5$\pm$0.8       64.3$\pm$2.2       47.6$\pm$0.8       33.9$\pm$2.8       61.6
      GroupDRO [14]       76.7$\pm$0.6       84.4$\pm$0.8       66.0$\pm$0.7       43.2$\pm$1.1       33.3$\pm$0.2       60.7
      Mixup [20]       77.4$\pm$0.6       84.6$\pm$0.6       68.1$\pm$0.3       47.9$\pm$0.8       39.2$\pm$0.1       63.4
      MLDG [15]       77.2$\pm$0.4       84.9$\pm$1.0       66.8$\pm$0.6       47.7$\pm$0.9       41.2$\pm$0.1       63.6
      CORAL [48]       78.8$\pm$0.6       86.2$\pm$0.3       68.7$\pm$0.3       47.6$\pm$1.0       41.5$\pm$0.1       64.6
      MMD [41]       77.5$\pm$0.9       84.6$\pm$0.5       66.3$\pm$0.1       42.2$\pm$1.6       23.4$\pm$9.5       58.8
      DANN [12]       78.6$\pm$0.4       83.6$\pm$0.4       65.9$\pm$0.6       46.7$\pm$0.5       38.3$\pm$0.1       62.6
      CDANN [42]       77.5$\pm$0.1       82.6$\pm$0.9       65.8$\pm$1.3       45.8$\pm$1.6       38.3$\pm$0.3       62.0
      MTL [44]       77.2$\pm$0.4       84.6$\pm$0.5       66.4$\pm$0.5       45.6$\pm$1.2       40.6$\pm$0.1       62.9
      SagNet [53]       77.8$\pm$0.5       86.3$\pm$0.2       68.1$\pm$0.1       48.6$\pm$1.0       40.3$\pm$0.1       64.2
      ARM [46]       77.6$\pm$0.3       85.1$\pm$0.4       64.8$\pm$0.3       45.5$\pm$0.3       35.5$\pm$0.2       61.7
      VREx [47]       78.3$\pm$0.2       84.9$\pm$0.6       66.4$\pm$0.6       46.4$\pm$0.6       33.6$\pm$2.9       61.9
      RSC [43]       77.1$\pm$0.5       85.2$\pm$0.9       65.5$\pm$0.9       46.6$\pm$1.0       38.9$\pm$0.5       62.7
      Mixstyle [100]       77.9$\pm$0.5       85.2$\pm$0.3       60.4$\pm$0.3       44.0$\pm$0.7       34.0$\pm$0.1       60.3
      AndMask [56]       78.1$\pm$0.9       84.4$\pm$0.9       65.6$\pm$0.4       44.6$\pm$0.3       37.2$\pm$0.6       62.0
      Fish [51]       77.8$\pm$0.3       85.5$\pm$0.3       68.6$\pm$0.4       45.1$\pm$1.3       42.7$\pm$0.2       63.9
      SelfReg [101]       77.8$\pm$0.9       85.6$\pm$0.4       67.9$\pm$0.7       47.0$\pm$0.3       42.8$\pm$0.0       64.2
      mDSDI [102]       79.0$\pm$0.3       86.2$\pm$0.2       69.2$\pm$0.4       48.1$\pm$1.4       42.8$\pm$0.1       65.1
      MIRO [64]       79.0$\pm$0.0       85.4$\pm$0.4       70.5$\pm$0.4       50.4$\pm$1.1       44.3$\pm$0.2       65.9
      SAM [34]       79.4$\pm$0.1       85.8$\pm$0.2       69.6$\pm$0.1       43.3$\pm$0.7       44.3$\pm$0.0       64.5
      SFT (Ours)       79.8$\pm$0.1       88.3$\pm$0.3       70.9$\pm$0.1       50.7$\pm$0.4       46.0$\pm$0.0       67.1
Table 2: DG performances on large-scale Vision Transformers. Out-of-domain accuracies of two backbones (ViT-B/16 and ViT-L/14) are shown below. The presence or absence of “*” indicates whether full fine-tuning or visual prompt tuning is employed, respectively.
     Backbone      Algorithms      VLCS      PACS      OfficeHome      TerraIncognita      DomainNet      Avg.
     ViT-B/16      ERM [45]      81.4      92.9      78.9      53.6      56.1      72.6
     MIRO  [64]      82.2      95.6      82.5      54.3      54.0      73.7
     ERM [45]      80.9      96.6      84.1      55.5      59.2      75.3
     IRM [50]      81.9      96.4      83.1      50.9      59.1      74.3
     DANN [12]      81.7      95.5      82.7      52.0      58.6      74.1
     CDANN [42]      81.9      96.0      82.3      54.9      58.4      74.7
     CORAL [48]      82.5      95.4      83.3      52.0      59.5      74.5
     MMD [41]      81.9      95.1      83.7      56.9      59.9      75.5
     IIB [22]      82.5      96.0      83.9      58.0      58.6      75.8
     SAM [34]      83.5      96.1      85.7      56.6      59.8      76.3
     SFT (Ours)      84.1      96.8      86.5      61.2      60.5      77.8
     ViT-L/14      ERM [45]      82.9      98.8      90.2      61.3      65.4      79.7
     SAM [34]      84.2      98.7      90.8      62.8      65.2      80.3
     SFT (Ours)      84.4      98.6      91.3      65.2      66.5      81.2
4.2.1 Experimental setting
Dataset and protocol. We evaluate our method on five widely-used and challenging datasets for domain generalization, including PACS [78] (4 domains, 7 classes, and 9, 991 examples), VLCS [79] (4 domains, 5 classes, and 10, 729 examples), OfficeHome [80] (4 domains, 65 classes, and 15, 588 images), TerraIncognita [81] (4 domains, 10 classes, and 24, 788 examples) and DomainNet [82] (6 domains, 345 classes, and 586, 575 examples). Our experimental environment is built based on the well-known DomainBed benchmark [11]. Following the training and evaluation protocol of DomainBed, we select one domain for testing and the remaining domains for training every time, and 20% samples of training domains are held out for validation and model selection. Model selection is carried out based on the training-domain validation set.
Implementation details. To illustrate the broad applicability of the proposed method, we conduct experiments based on small models (ImageNet pre-trained ResNet-50) and large-scale pre-trained models (ViT-B/16 and ViT-L/14 for CLIP  [69]). Unless otherwise specified, the landscape refiner shares the same architecture with the model $\boldsymbol{f}_{\boldsymbol{\theta}}$ in this paper. Through random hyperparameter search, we determine the batch size, learning rates, and other hyperparameters. The detailed search spaces for each are provided in the supplementary file. The Adam [23] optimizer is utilized in all of our experiments.
4.2.2 Main Results
Experiments with the ResNet-50 backbone. We begin by comparing SFT with conventional DG methods. As shown in Table 1, SFT outperforms ERM [45] across all five datasets, showing an average improvement of 3.8%. Furthermore, SFT consistently outperforms other conventional DG methods, including invariant or causal representation learning methods (IRM [50], VREx [47], CORAL [48], MMD [41], DANN [12], CDANN [42]), disentangled representation learning methods (SagNet [53] and mDSDI [102]), distributionally robust optimization methods (GroupDRO [14]), meta-learning methods (MLDG [15], ARM [46], and Fish [51]), data or feature augmentation methods (Mixup [20] and Mixstyle [100]), and other methods (RSC [43], AndMask [56], MTL [44], and SelfReg [101]). As a strong baseline, SAM achieves an average improvement of +1.2% (63.3% $\rightarrow$ 64.5%) over ERM and outperforms most other baselines, except for mDSDI [102] and MIRO [64]. By refining loss landscapes to improve consistency, SFT provides an additional average gain of +2.6% (64.5% $\rightarrow$ 67.1%) over SAM. Remarkably, our SFT even outperforms the robust MIRO [64] baseline by 1.2%.
Experiments with Large-scale Vision Transformers (ViTs). To demonstrate the versatility of SFT across different architectures, we also conduct experiments using large-scale ViTs. Given the efficiency of Visual Prompt Tuning (VPT) [66]—which requires tuning only 1% of the parameters—and its strong performance in domain generalization, as shown in Table 2 and previous works [67, 68], we adopt VPT in our ViT experiments. From Table 2, we can see that SFT achieves superior performance over conventional DG methods and the vanilla SAM (+1.5% average improvement) when using ViT-B/16. When using a larger pre-trained ViT-L/14, the out-of-domain accuracies further improve, achieving 79.7% for ERM, 80.3% for SAM, and 81.2% for SFT. The results highlight that SFT can be easily integrated with popular tuning methods like VPT, and there is potential for further enhancement with advanced tuning techniques in future work.
Comparison with popular sharpness-aware methods. We also perform a more fine-grained comparison of SFT with popular SAM variants designed to seek better flat minima. As shown in Table 3, SFT generally outperforms these methods across all five datasets, with average improvements of 2.7%, 2.0%, 1.8% and 1.0% over GAM [103], GSAM [96], FAD [104] and SAGM [95], respectively. These experimental results further highlight the importance of the consistency of flat minima. Since SFT currently only applies the basic SAM algorithm, there is potential for further improvements by incorporating more advanced SAM variants, such as SAGM [95], which remains a focus for our future research.
Table 3: Comparisons of SFT and popular sharpness-aware methods. This table presents average DG performances on five datasets, and the mean and standard deviation for each dataset.
Algorithms VLCS PACS OfficeHome TerraIncognita DomainNet Avg.
SAM [34] 79.4$\pm$0.1 85.8$\pm$0.2 69.6$\pm$0.1 43.3$\pm$0.7 44.3$\pm$0.0 64.5
GAM [103] 78.5$\pm$0.4 86.1$\pm$0.6 68.2$\pm$1.0 45.2$\pm$0.6 43.8$\pm$0.1 64.4
GSAM [96] 79.1$\pm$0.2 85.9$\pm$0.1 69.3$\pm$0.0 47.0$\pm$0.8 44.6$\pm$0.2 65.1
FAD [104] 78.9$\pm$0.8 88.2$\pm$0.5 69.2$\pm$0.5 45.7$\pm$1.0 44.4$\pm$0.1 65.3
SAGM [95] 80.0$\pm$0.3 86.6$\pm$0.2 70.1$\pm$0.2 48.8$\pm$0.9 45.0$\pm$0.2 66.1
SFT (Ours) 79.8$\pm$0.1 88.3$\pm$0.3 70.9$\pm$0.1 50.7$\pm$0.4 46.0$\pm$0.0 67.1
Comparison with weight average methods. Weight averaging methods, such as SWAD [70], seek flat minima in a different manner. Previous work [72] has compared it with SAM and revealed that parameters sought by the two methods are located in different basins. Since model averaging is a totally different method of model selection, it wouldn’t be fair to make a direct comparison with SFT. Nonetheless, in our experiment with ViT-B/16 (as shown in Table 4), we observed that SAM and SFT outperform SWAD by 0.6% (85.1$\rightarrow$85.7%) and 1.4% (85.1$\rightarrow$86.5%), respectively.
4.3 Ablation Study
Ablation studies on the key components of the SFT framework are summarized in Table 5. The upper part of the table shows experimental results for ERM and SAM with one-hot and smoothed labels, while the lower part presents results using soft labels generated by the landscape refiner. The results in the upper part confirm the effectiveness of SAM (where $\rho\neq 0$) and label smoothing. When comparing the six experiments in the lower part, we observe that incorporating all components of the SFT framework to enhance landscape consistency across domains yields the best performance, achieving an average accuracy of 86.5%. Furthermore, removing any component from the SFT framework results in a performance drop, which underscores the importance of each element within the framework. Finally, the comparison between the upper and lower sections demonstrates that the consistency-guided soft labels in our SFT framework offer a clear advantage over traditional one-hot or smoothed labels, significantly improving model generalization.
Table 4: Comparisons of SFT and the weight average method (SWAD). We report out-of-domain accuracies with/without SWAD on four domains of OfficeHome.
Algorithms Art Clipart Product Real World Avg.
ERM [45] 83.7 73.8 89.9 89.2 84.1
ERM [45]$+$SWAD [70] 84.9 74.9 90.8 89.9 85.1
SAM [34] 84.2 76.6 91.0 90.8 85.7
SAM [34]$+$SWAD [70] 84.6 76.5 91.2 90.8 85.8
SFT (Ours) 86.3 77.7 91.1 90.9 86.5
SFT (Ours) $+$ SWAD [70] 86.0 77.5 91.2 91.3 86.5
Table 5: Ablative experiments on the OfficeHome dataset. The “Label” column specifies the type of labels used, such as one-hot labels, smoothed labels or labels generated by the landscape refiner. The “PCE” indicates whether the PCE loss is utilized in the Exp.5-10. If not, we use the standard cross entropy with label smoothing instead.
Exp. $\rho$ Label PCE $\lambda_{1}$ $\lambda_{2}$ Art Clipart Product Real World Avg.
1 0.0 One-hot -- -- -- 83.2 73.8 90.1 89.1 84.0
2 0.0 smoothed -- -- -- 83.6 74.1 90.0 89.6 84.3
3 0.3 One-hot -- -- -- 83.8 76.6 90.1 90.3 85.2
4 0.3 Smoothed -- -- -- 84.2 76.4 90.6 90.5 85.4
5 0.0 Generated w/o 0.0 0.0 83.9 74.9 89.9 90.0 84.7
6 0.3 Generated w/o 0.0 0.0 84.0 75.9 90.1 90.1 85.0
7 0.3 Generated w/o 0.5 0.0 84.8 76.8 90.3 90.6 85.6
8 0.3 Generated w/o 0.0 0.7 84.9 77.3 90.6 90.5 85.8
9 0.3 Generated w/o 0.5 0.7 85.2 77.2 90.8 90.7 86.0
10 0.3 Generated w/ 0.5 0.7 86.3 77.7 91.2 90.9 86.5
5 Conclusion
This paper has proposed an iterative two-phase Self-Feedback Training (SFT) framework, aiming at addressing the issue of landscape inconsistency brought by domain shifts. It alternatively generates a feedback signal by measuring the inconsistency of loss landscapes in different domains during the feedback phase, and then refines these loss landscapes for greater consistency using this feedback signal in the refinement phase. Benefiting from the consistency of the sought flat minima, SFT demonstrates superior performances over sharpness-aware methods and prevalent DG methods in diverse experimental settings. We hope that this study will inspire more research on advanced DG methods from the view of loss landscapes. Further theoretical analysis and technical improvement will be our future work.
Supplementary Material
Let us start with a brief overview of this supplementary material. In Section 6, we perform a theoretical analysis of the loss landscape consistency under the PAC-Bayesian framework. The analysis shows that the sharpness of the test loss can be constrained by that of the training domains, which also implies the generalization of consistent flat minima sought by the self-feedback training (SFT). In Section 7, we discuss the proposed projection cross-entropy loss and the algorithm used to solve the associated KL divergence minimization problem. We provide a comprehensive derivation of the algorithm, along with a time comparison against an implementation using the common convex optimization library. Subsequently, we present the full results obtained using Resnet-50, ViT-B/16, and ViT-L/14 in the following two sections (Section 8 and 9). Additionally, we offer necessary explanations of the code (in Section 10.1) and hyperparameters (in Section 10.2) required for reproducibility. Finally, in Section 11, we assure that this research is unlikely to have any significant negative social impact.
6 Theoretical Analysis
In this section, we will perform a theoretical analysis of the loss landscape consistency under the PAC-Bayesian framework. For clarity and ease of understanding, we first provide a detailed explanation of the relevant notations and concepts that will be used throughout the analysis.
6.1 Notations
Let $\mathcal{X}$ and $\mathcal{Y}$ denote the input sample space and the category space, respectively. Consider a dataset drawn from $p$ training distributions $\{\mathcal{D}_{d}\}_{d=1}^{p}$, each defined over the joint space $\mathcal{X}\times\mathcal{Y}$. Let $D_{d}=\{(\boldsymbol{x}_{i}^{(d)},y_{i}^{(d)})\}_{i=1}^{n_{d}}$ denote the dataset sampled from the $d$-th distribution $\mathcal{D}_{d}$, which is referred to as the $d$-th domain. $(\boldsymbol{x}_{i}^{(d)},y_{i}^{(d)})\in\mathcal{X}\times\mathcal{Y}$ denotes the $i$-th sample from domain $\mathcal{D}_{d}$, with $n_{d}$ indicating the number of samples in the $d$-th domain. For convenience, we also use $z_{i}$ to denote $(\boldsymbol{x}_{i},y_{i})$. Let $\Omega$ and $\Omega^{\prime}$ denote the dataset space and distribution space, respectively. In our analysis, domain shifts are modeled by a mapping function $\omega:\Omega\rightarrow\Omega^{\prime}$, which maps one dataset to another distribution with distinct statistical properties. We assume that the domain shifts $\omega$ follow a specific distribution $\mathcal{W}$.
Our SFT framework mainly involves a model and a landscape refiner. Let $\mathcal{H}_{m}$ and $\mathcal{H}_{r}$ denote the hypothesis spaces of the model and the refiner, respectively. To analyze the SFT within the PAC-Bayesian framework, we need to provide a description using Bayesian terms. Let $\mathcal{M}_{m}$, $\mathcal{M}_{r}$ denote the sets of distributions over $\mathcal{H}_{m}$ and $\mathcal{H}_{r}$. In order to obtain a model from its prior distribution, we first sample a prior distribution $P$ for the refiner from the hyper-prior distribution $\mathcal{P}$, which is independent of training samples. Then, we use a mapping function $\psi:\mathcal{M}_{r}\rightarrow\mathcal{M}_{m}$ to obtain the model’s prior distribution $\psi(P)\in\mathcal{M}_{m}$ and sample a model $f$ from $\psi(P)$. In order to obtain a model from its posterior distribution, we sample a posterior distribution $P$ for the refiner from a hyper-posterior distribution $\mathcal{Q}$, which may depend on training samples. Then, we apply the training algorithm to the dataset $D_{d}\in\Omega$ to obtain the model’s posterior distribution $\mathcal{A}(D_{d},P)$, where $\mathcal{A}:\Omega\times\mathcal{M}_{r}\rightarrow\mathcal{M}_{m}$ represents the function that maps one dataset along with the refiner’s posterior distribution to the model’s posterior distribution. Finally, we can sample a model $f$ from this posterior distribution.
6.2 Main Theorem
In this subsection, we first introduce a lemma that will be used multiple times in our analysis, and then formally present the main result, which is stated as Theorem 2.
Lemma 1 (McAllester’s bound [105]).
Let $\mathcal{X}$ be a sample space and $\mathcal{H}$ a hypothesis space of functions over $\mathcal{X}$. Given $\pi$ be some prior distribution over hypothesis space $\mathcal{H}$, for bounded loss :×$\ell:\mathcal{H}\times\mathcal{X}\rightarrow[0,1]$ and any $\delta\in(0,1]$, the following bound holds uniformly for all posterior distributions $\rho$ with probability at least $1-\delta$:
$\underset{\theta\sim\rho}{\mathbb{E}}\ell(\theta,\mathcal{D})\leq\underset{\theta\sim\rho}{\mathbb{E}}\ell(\theta,S_{n})+\sqrt{\frac{\mathrm{KL}(\rho||\pi)+\log{(n/\delta)}}{2(n-1)}},$ (14)
where
$\ell(\theta,\mathcal{D})=\mathbb{E}_{z\sim D}\ell(\theta,z)$
and
$\ell(\theta,S_{n})=\frac{1}{n}\sum_{i=1}^{n}\ell(\theta,z_{i})$
denotes the population loss and training loss, respectively. $S_{n}$ represents a dataset with $n$ training samples drawn independently and identically from distribution $\mathcal{D}$.
Theorem 2.
Consider the domain generalization problem with $p$ training domains. For each training domain $\mathcal{D}_{d}$, we are given $p-1$ training-domain pairs $(\mathcal{D}_{d},\mathcal{D}_{d^{\prime}})$, where $d^{\prime}\neq d$, with each pair consisting of a training dataset $D_{d}$ of size $n_{d}$ and a hold-out dataset $D_{d^{\prime}}$ of size $n_{d^{\prime}}$. Let the dataset space be $\Omega$ and the distribution space be $\Omega^{\prime}$. Assume that the domain shifts, denoted by $\omega:\Omega\rightarrow\Omega^{\prime}$, follow a distribution $\mathcal{W}$. The difference in loss sharpness between domains $\mathcal{D}_{d}$ and $\mathcal{D}_{d^{\prime}}$ is defined as:
$\Delta\ell(f,\mathcal{D}_{d},\mathcal{D}_{d^{\prime}}):=|\underset{z\sim\mathcal{D}_{d}}{\mathbb{E}}\ell(f,z)-\underset{z\sim\mathcal{D}_{d^{\prime}}}{\mathbb{E}}\ell(f,z)|,$ (15)
where $\ell(f,z)$ represent the loss sharpness of the model $f$ evaluated on data point $z$. Let $\mathcal{P}$ denote a predefined hyper-prior distribution over the set of all possible prior distributions for the landscape refiner. Then, for all hyper-posterior distributions $\mathcal{Q}$ that ensure a sufficiently small sharpness difference between training domains, i.e.,
$\forall d^{\prime}\neq d:\underset{P\sim\mathcal{Q}}{\mathbb{E}}\ \underset{f\sim\mathcal{A}(D_{d},P)}{\mathbb{E}}\Delta\ell(f,\mathcal{D}_{d},\mathcal{D}_{d^{\prime}})\leq\epsilon,$ (16)
and for any $\delta\in(0,1]$, the following inequality holds with probability at least $1-\delta$:
$\quad\ell(\mathcal{Q},\omega)\leq\epsilon+\hat{\ell}(\mathcal{Q},D_{d})+\frac{1}{p-1}\times$
$\sum_{d^{\prime}\neq d}\sqrt{\frac{\mathrm{KL}(\mathcal{Q}\|\mathcal{P})+\underset{P\sim\mathcal{Q}}{\mathbb{E}}\mathrm{KL}\left(\mathcal{A}\|\psi(P)\right)+\log\frac{2(p-1)n_{d}}{\delta}}{2(n_{d}-1)}}$
$+\sqrt{\frac{\mathrm{KL}(\mathcal{Q}\|\mathcal{P})+\log\frac{2(p-1)}{\delta}}{2(p-2)}}.$ (17)
Here, $\psi(P)$ is the prior model distribution, and $\mathcal{A}$ is a shorthand of the posterior model distribution $\mathcal{A}(D_{d},P)$. $\mathrm{KL}(\cdot\|\cdot)$ denotes the Kullback-Leibler divergence. The term $\ell(\mathcal{Q},\omega)$ is defined as:
$\ell(\mathcal{Q},\omega):=\underset{P\sim\mathcal{Q}}{\mathbb{E}}\ \underset{\omega\sim\mathcal{W}}{\mathbb{E}}\ \underset{f\sim\mathcal{A}(D_{d},P)}{\mathbb{E}}\ \underset{z\sim\omega(D_{d})}{\mathbb{E}}\ell(f,z),$ (18)
which represents the expected sharpness of the model evaluated on the test distribution $\omega(D_{d})$. The term ^$\hat{\ell}(\mathcal{Q},D_{d})$ is defined as:
^$\hat{\ell}(\mathcal{Q},D_{d}):=\underset{P\sim\mathcal{Q}}{\mathbb{E}}\ \underset{f\sim\mathcal{A}(D_{d},P)}{\mathbb{E}}\frac{1}{n_{d}}\sum_{i=1}^{n_{d}}\ell\left(f,z_{i}\right),$ (19)
which represents the empirical sharpness of the model over the training domain $D_{d}$.
Proof.
Firstly, we bound the loss sharpness in each of the domain pairs. Based on the above definitions in the subsection 6.1, we can decompose the KL divergence term of Lemma 1 in the following way:
$\quad\mathrm{KL}(\rho\|\pi)$
$=\underset{f\sim\rho}{\mathbb{E}}\log\frac{\rho(f)}{\pi(f)}=\underset{P\sim\mathcal{Q}}{\mathbb{E}}\ \underset{f\sim\mathcal{A}(D_{d},P)}{\mathbb{E}}\log\frac{\mathcal{Q}(P)\mathcal{A}\left(D_{d},P\right)(f)}{\mathcal{P}(P)\psi(P)(f)}$
$=\underset{P\sim\mathcal{Q}}{\mathbb{E}}\log\frac{\mathcal{Q}(P)}{\mathcal{P}(P)}+\underset{P\sim\mathcal{Q}}{\mathbb{E}}\ \underset{f\sim\mathcal{A}(D_{d},P)}{\mathbb{E}}\log\frac{\mathcal{A}\left(D_{d},P\right)(f)}{\psi(P)(f)}$
$=\mathrm{KL}(\mathcal{Q}\|\mathcal{P})+\underset{P\sim\mathcal{Q}}{\mathbb{E}}\mathrm{KL}\left(\mathcal{A}\left(D_{d},P\right)\|\psi(P)\right).$ (20)
This decomposition separates the KL divergence into two components: $\mathrm{KL}(\mathcal{Q}\|\mathcal{P})$ and $\underset{P\sim\mathcal{Q}}{\mathbb{E}}\mathrm{KL}\left(\mathcal{A}\left(D_{d},P\right)\|\psi(P)\right)$. Now, based on this, we can establish a probabilistic upper bound on the loss sharpness. Thus, with probability at least $1-\delta_{d^{\prime}}$, we have:
$\quad\underset{P\sim\mathcal{Q}}{\mathbb{E}}\ \underset{f\sim\mathcal{A}\left(D_{d},P\right)}{\mathbb{E}}\ \underset{z\sim\mathcal{D}_{d^{\prime}}}{\mathbb{E}}\ell(f,z)$
$\leq\underset{P\sim\mathcal{Q}}{\mathbb{E}}\ \underset{f\sim\mathcal{A}\left(D_{d},P\right)}{\mathbb{E}}\frac{1}{n_{d}}\sum_{i=1}^{n_{d}}\ell\left(h,z_{i}\right)$
$\quad+\underset{P\sim\mathcal{Q}}{\mathbb{E}}\ \underset{f\sim\mathcal{A}\left(D_{d},P\right)}{\mathbb{E}}\Delta\ell(h,\mathcal{D}_{d},\mathcal{D}_{d^{\prime}})$
$\quad+\sqrt{\frac{\mathrm{KL}(\mathcal{Q}\|\mathcal{P})+\underset{P\sim\mathcal{Q}}{\mathbb{E}}\mathrm{KL}\left(\mathcal{A}\|\psi(P)\right)+\log\frac{n_{d}}{\delta_{d^{\prime}}}}{2\left(n_{d}-1\right)}}.$ (21)
This bound captures the generalization ability between training domains, i.e., from domain $\mathcal{D}_{d}$ to domain $\mathcal{D}_{d^{\prime}}$.
Next, assuming that the domain shifts $\omega:\Omega\rightarrow\Omega^{\prime}$ are governed by a distribution $\mathcal{W}$, we can apply Lemma 1 again to bound the loss sharpness on the test domain $\omega(D_{d})$. This step is critical because it extends the generalization to the test domains. Specifically, with probability at least $1-\delta^{\prime}$, we obtain the following bound:
$\quad\underset{P\sim\mathcal{Q}}{\mathbb{E}}\ \underset{\omega\sim\mathcal{W}}{\mathbb{E}}\ \underset{f\sim\mathcal{A}(D_{d},P)}{\mathbb{E}}\ \underset{z\sim\mathcal{\omega}(D_{d})}{\mathbb{E}}\ell(f,z)$
$\leq\underset{P\sim\mathcal{Q}}{\mathbb{E}}\frac{1}{p-1}\sum_{d^{\prime}\neq d}\underset{f\sim\mathcal{A}(D_{d},P)}{\mathbb{E}}\ \underset{z\sim\mathcal{D}_{d^{\prime}}}{\mathbb{E}}\ell(f,z)$
$+\sqrt{\frac{\mathrm{KL}(\mathcal{Q}\|\mathcal{P})+\log\frac{p-1}{\delta^{\prime}}}{2(p-2)}}.$ (22)
Here, the first term represents an average over all domain pairs, while the second term corresponds to the KL divergence between hyper-prior and hyper-posterior distributions. The confidence parameter $\delta^{\prime}$ controls the probability of the bound holding.
Finally, we set $\delta^{\prime}=\delta/2$ and $\delta_{d^{\prime}}=\delta/[2(p-1)]$ and apply the union bound to combine the above results. This leads to the final bound, which holds with probability at least $1-\delta$,
$\quad\underset{P\sim\mathcal{Q}}{\mathbb{E}}\ \underset{\omega\sim\mathcal{W}}{\mathbb{E}}\ \underset{f\sim\mathcal{A}(D_{d},P)}{\mathbb{E}}\ \underset{z\sim\omega(D_{d})}{\mathbb{E}}\ell(h,z)$
$\leq\frac{1}{p-1}\sum_{d^{\prime}\neq d}\left[\underset{P\sim\mathcal{Q}}{\mathbb{E}}\ \underset{f\sim\mathcal{A}\left(D_{d},P\right)}{\mathbb{E}}\frac{1}{n_{d}}\sum_{i=1}^{n_{d}}\ell\left(f,z_{i}\right)\right.$
$\left.+\underset{P\sim\mathcal{Q}}{\mathbb{E}}\ \underset{f\sim\mathcal{A}\left(D_{d},P\right)}{\mathbb{E}}\Delta\ell(f,\mathcal{D}_{d},\mathcal{D}_{d^{\prime}})\right.+$
$\left.\sqrt{\frac{\mathrm{KL}(\mathcal{Q}\|\mathcal{P})+\underset{P\sim\mathcal{Q}}{\mathbb{E}}\mathrm{KL}\left(\mathcal{A}\|\psi(P)\right)+\log\frac{2(p-1)n_{d}}{\delta}}{2(n_{d}-1)}}\right]$
$+\sqrt{\frac{\mathrm{KL}(\mathcal{Q}\|\mathcal{P})+\log\frac{2(p-1)}{\delta}}{2(p-2)}}.$ (23)
The inequality above provides a comprehensive bound of the loss sharpness on the test domains, completing the proof of Theorem 2. ∎
In conclusion, this theorem shows that if domain shifts are governed by a specific distribution and the domain shifts in the training domains are independently and identically sampled from this distribution, then the sharpness of the test loss is bounded by the sharpness observed in the training domains with high probability. In other words, the consistency of the flat minima achieved in the training domains can be transferred to unseen test domains. As a result, the model can exhibit strong generalization performance when applied to test domains.
7 Projection Cross Entropy
As mentioned in the main text, the projection cross entropy (PCE) can be used as a loss term to maintain the label correctness during the refinement phase. An efficient algorithm (Alogrithm 2) has been presented to address the associated KL divergence minimization problem there.
In this section, we first provide the derivation of this algorithm, followed by a comparison of its time efficiency with that of widely-used convex optimization libraries. Finally, we further discuss the connections between the PCE loss and other related loss functions.
7.1 KL Divergence Minimization
As stated in the main text, the optimization is formulated as:
$\min_{\mathbf{y}}\,\mathrm{KL}(\mathbf{y}\,||\,\tilde{\mathbf{y}})\quad\text{subject to}\quad\mathbf{y}\in C_{1},$ (24)
where $\tilde{\textbf{y}}$ represents the soft label output by the landscape refiner, and the label space $C_{1}$ can be expressed as:
$C_{1}=\{(q_{1},\ldots,q_{N})\mid\forall k\neq 1:q_{1}\geq\alpha q_{k},\sum_{k=1}^{N}q_{k}=1\}.$ (25)
Here, $\alpha\geq 1$ is a hyperparameter that controls the minimum ratio between $q_{1}$ and $q_{k}$, while $N$ denotes the number of categories for classification. For ease of understanding, we restate the problem more explicitly as follows:
$\min\limits_{q_{i}}\sum_{i=1}^{N}q_{i}\log\frac{q_{i}}{p_{i}}$
s. t. $\sum_{i=1}^{N}q_{i}=1,\quad q_{1}\geq\alpha q_{k}\ (k\neq 1),$ (26)
where we use $(p_{1},\ldots,p_{N})$ to represent $\tilde{\textbf{y}}$ for clarity.
In the following, we will offer a detailed derivation of Algorithm 2, which is capable of finding the exact optimal solution to this optimization problem. The general idea of the derivation is as follows: first, by examining the Lagrangian function and the Karush-Kuhn-Tucker (KKT) conditions of the problem, we obtain the general form of the optimal solution. Then, we determine which constraints among the inequality constraints hold as a strict equality (i.e., active), and finally, we find the optimal solution based on these active constraints.
7.1.1 Lagrangian Function and KKT Conditions
To solve the given optimization problem, we first construct the Lagrangian function by incorporating the objective function and the constraints using Lagrange multipliers:
$L(q_{i},\mu_{k},\lambda)$ $=\sum_{i=1}^{N}q_{i}\log\frac{q_{i}}{p_{i}}+\sum_{k=2}^{N}\mu_{k}(\alpha q_{k}-q_{1})$
$+\lambda(\sum_{i=1}^{N}q_{i}-1),$ (27)
where $\mu_{k}\geq 0$ represents the Lagrange multiplier for the inequality constraint $q_{1}\geq\alpha q_{k}$, and $\lambda$ denotes the Lagrange multiplier for the equality constraint $\sum_{i=1}^{N}q_{i}=1$. Then, the Karush-Kuhn-Tucker (KKT) conditions are the necessary conditions for optimality in constrained optimization problems. We apply these conditions to the Lagrangian function and derive the following set of equations:
  1. 1.
    Stationarity.  The stationarity conditions require that the partial derivatives of the Lagrangian with respect to each of the variables be zero, which corresponds to the optimality condition. For $q_{1}$, we have the following equation:
    $\frac{\partial L}{\partial q_{1}}=1+\log{q_{1}}-\log{p_{1}}-\sum_{k=2}^{N}\mu_{k}+\lambda=0.$ (28)
    Similarly, for $q_{j}$, where $j=2,\ldots,n$, we get:
    $\frac{\partial L}{\partial q_{j}}=1+\log{q_{j}}-\log{p_{j}}+\mu_{j}\alpha+\lambda=0.$ (29)
    These two equations can be solved to express $q_{1}$ and $q_{j}$ in terms of the Lagrange multipliers $\mu_{k}$ and $\lambda$. Thus, we have the following solutions:
    $q_{1}=p_{1}\exp{\left(\sum_{k=2}^{N}\mu_{k}-1-\lambda\right)}$ (30)
    and
    $q_{j}=p_{j}\exp{\left(-\mu_{j}\alpha-1-\lambda\right)}.$ (31)
  2. 2.
    Primal feasibility. The primal feasibility condition ensures that the original constraints are satisfied. Therefore, we have:
    $\alpha q_{j}-q_{1}\leq 0.$ (32)
  3. 3.
    Dual feasibility. The dual feasibility condition imposes non-negativity on the Lagrange multipliers associated with the inequality constraints:
    $\mu_{j}\geq 0.$ (33)
  4. 4.
    Complementary slackness. Finally, the complementary slackness condition relates the primal and dual variables. In this case, the complementary slackness condition for the inequality constraint is:
    $\mu_{j}(\alpha q_{j}-q_{1})=0.$ (34)
    This condition implies that either $\mu_{j}=0$ or $\alpha q_{j}=q_{1}$ (or both). It ensures that if the constraint is inactive (i.e., it holds as a strict inequality), the corresponding multiplier is zero, and if the multiplier is positive, the corresponding constraint is active (i.e., it holds as a strict equality).
In the following, we can obtain the general form of the optimal solution by examining the two cases of $\mu>0$ and $\mu=0$ as per the KKT conditions.
  • Case 1: $\mu_{j}>0$.
    If $\mu_{j}>0$, from the stationarity condition and complementary slackness, we can derive the following equation:
    $\alpha q_{j}=\alpha p_{j}\exp{(-\mu_{j}\alpha-1-\lambda)}$
    $=$ $q_{1}=p_{1}\exp{\left(\sum_{k=2}^{N}\mu_{k}-1-\lambda\right)}.$ (35)
    Through a simple manipulation of the equation, we can obtain:
    $\frac{p_{1}}{\alpha p_{j}}=\frac{\exp(-\mu_{j}\alpha)}{\exp(\sum_{k=2}^{N}\mu_{k})}.$ (36)
    Since $\mu_{j}>0$ in this case, the right side of the above equation is less than 1. In other words, if $p_{1}\geq\alpha p_{j}$, then we can determine that $\mu_{j}=0$. On the other hand, by setting the value of $j$ in equation (7.1.1) and multiplying these expressions together, we can obtain the following relationship:
    $\alpha^{|A|}\exp(-\alpha\sum_{j\in A}\mu_{j})\prod_{j\in A}p_{j}=p_{1}^{|A|}\exp(|A|\sum_{k\in A}\mu_{k}).$ (37)
    Here, $A$ is defined as the set of indices $k$ such that $\mu_{k}>0$, i.e., $A:=\{k|\mu_{k}>0\}$. $|A|$ denotes the cardinality of set $A$. Then, we can derive:
    $\exp(\sum_{k\in A}\mu_{k})=(p_{1}^{-|A|}(\prod_{j\in A}\alpha p_{j}))^{\frac{1}{|A|+\alpha}}.$ (38)
    By using the above equation, we can express $q_{1}$ and $q_{j}$ without using $\mu_{j}$:
    $q_{1}=\alpha q_{j}=\exp{(-1-\lambda)}(p_{1}^{\alpha}(\prod_{j\in A}\alpha p_{j}))^{\frac{1}{|A|+\alpha}}.$ (39)
  • Case 2: $\mu_{j}=0$.
    If $\mu_{j}=0$, $q_{j}$ can be easily expressed without using $\mu_{j}$:
    $q_{j}=p_{j}\exp{(-1-\lambda)}.$ (40)
    Then, by using the equations (7.1.1), (39) and (40), we can also express the objective function without using $\mu_{j}$:
    $\sum_{i=1}^{N}q_{i}\log\frac{q_{i}}{p_{i}}=-(1+\lambda).$ (41)
    Finally, by applying the normalization condition $\sum_{i=1}^{N}q_{i}=1$, we can solve for the value of $\lambda$ using the following equation:
    $\exp(1+\lambda)=(1+\frac{|A|}{\alpha})(p_{1}^{\alpha}(\prod_{j\in A}\alpha p_{j}))^{\frac{1}{|A|+\alpha}}+\sum_{j\not\in A\cup\{1\}}p_{j}.$ (42)
7.1.2 Determine Active Constraints
To determine the active constraints, we have the following proposition, which directly leads to the formulation of Algorithm 2.
Proposition 3.
Consider the optimization problem in (26). Let $A$ be the set of indices $k$ such that $\mu_{k}>0$, i.e., $A:=\{k|\mu_{k}>0\}$. Define $B$ as the set $\{p_{j}|\alpha p_{j}>p_{1}\}$. Sort the elements of $B$ in descending order as:
$p_{j_{1}}\geq p_{j_{2}}\geq\ldots\geq p_{j_{|B|}}.$ (43)
Note that duplicate elements in $B$ are not removed. Then, the following conclusions holds:
  1. 1.
    The index $j_{1}$, which corresponds to the largest element in $B$, must belong to $A$, i.e., $j_{1}\in A$.
  2. 2.
    For any $C=\{j_{1},\ldots,j_{t-1}\}\subseteq A$, if the inequality
    $(p_{1}^{\alpha}(\prod_{j\in C}\alpha p_{j}))^{\frac{1}{|C|+\alpha}}<\alpha p_{j_{t}}$ (44)
    holds, then $j_{t}\in A$. Otherwise, for all $s\in\{t,t+1,\ldots,|B|\}$, we have $j_{s}\not\in A$. That is, $A=C$.
Proof.
Firstly, we can prove that $j_{1}\in A$. To do so, assume for the sake of contradiction that $j_{1}\notin A$. By using equations (39) and (40), we obtain the following expression for $q_{1}$:
$q_{1}$ $=\exp{(-1-\lambda)}(p_{1}^{\alpha}(\prod_{j\in A}\alpha p_{j}))^{\frac{1}{|A|+\alpha}}$
$<\exp{(-1-\lambda)}((\alpha p_{j_{1}})^{\alpha}(\prod_{j\in A}\alpha p_{j_{1}}))^{\frac{1}{|A|+\alpha}}$
$=\alpha\exp{(-1-\lambda)}p_{j_{1}}=\alpha q_{j_{1}}.$ (45)
Here, the first equality follows from the equation (39), and the inequality in the second-to-last line is based on the definition of B, in which $p_{1}<\alpha p_{j_{1}}$. The first equality in the last line is simply an identity transformation, and the last equation uses the equation (40) under the assumption that $j_{1}\notin A$. The above analysis shows that $q_{1}<\alpha q_{j_{1}}$, which clearly violates the inequality constraints of the original problem. Thus, our assumption that $j_{1}\notin A$ must be false.
Secondly, we consider the case where $C=\{j_{1},\ldots,j_{t-1}\}\subseteq A$, and we assume that the inequality
$(p_{1}^{\alpha}(\prod_{j\in C}\alpha p_{j}))^{\frac{1}{|C|+\alpha}}<\alpha p_{j_{t}}$ (46)
holds. We now want to prove that $j_{t}\in A$. Suppose, for contradiction, that $j_{t}\notin A$. In this case, we proceed as follows. Using the equations (39) and (40), we can express $q_{1}$ as:
$q_{1}$ $=\exp{(-1-\lambda)}(p_{1}^{\alpha}(\prod_{j\in A}\alpha p_{j}))^{\frac{1}{|A|+\alpha}}$
$=\exp{(-1-\lambda)}(p_{1}^{\alpha}(\prod_{j\in C}\alpha p_{j})(\prod_{j\in A-C}\alpha p_{j}))^{\frac{1}{|A|+\alpha}}$
$<\exp{(-1-\lambda)}(p_{1}^{\alpha}(\prod_{j\in C}\alpha p_{j})(\prod_{j\in A-C}\alpha p_{j_{t}}))^{\frac{1}{|A|+\alpha}}$
$<\exp{(-1-\lambda)}((\alpha p_{j_{t}})^{|C|+\alpha}(\prod_{j\in A-C}\alpha p_{j_{t}}))^{\frac{1}{|A|+\alpha}}$
$=\alpha\exp{(-1-\lambda)}p_{j_{t}}=\alpha q_{j_{t}}.$ (47)
The first equality uses equation (39). The first inequality follows from the fact that $p_{j_{t}}>p_{j}$ for any $j\in A-C$, which is because the elements of $B$ have been sorted in descending order. The second inequality uses the condition stated in (46). The last equality follows from equation (40) under the assumption that $j_{t}\notin A$. The above analysis shows that $q_{1}<\alpha q_{j_{t}}$, which, again, violates the inequality constraints of the original problem. Hence, our assumption that $j_{t}\notin A$ is false, and it holds that $j_{t}\in A$.
Finally, we address the case where the inequality (46) does not holds. In this case, for all $s\in\{t,t+1,\ldots,|B|\}$, we assert that $j_{s}\not\in A$. We will now prove this by considering two cases. In the first case, we assume that there exists some $j_{s}\in A$ such that $p_{1}^{\alpha}(\prod_{j\in C}\alpha p_{j}))^{\frac{1}{|C|+\alpha}}>\alpha p_{j_{s}}$. If $j_{s}\in A$, we then have:
$\quad\sum_{i=1}^{N}q_{i}\log\frac{q_{i}}{p_{i}}=-(1+\lambda)$
$=-\log\left[(1+\frac{|C|+1}{\alpha})((p_{1}^{\alpha}(\prod_{j\in C\cup\{j_{s}\}}\alpha p_{j}))^{\frac{1}{|C|+1+\alpha}}\right.$
$\left.\quad+\sum_{j\not\in C\cup\{1,j_{s}\}}p_{j}\right].$ (48)
The equations (41) and (42) have been applied above. However, if $j_{s}\not\in A$, we have
$\quad\sum_{i=1}^{N}q_{i}\log\frac{q_{i}}{p_{i}}=-(1+\lambda)$
$=-\log\left[(1+\frac{|C|}{\alpha})(p_{1}^{\alpha}(\prod_{j\in C}\alpha p_{j}))^{\frac{1}{|C|+\alpha}}+\sum_{j\not\in C\cup\{1\}}p_{j}\right].$ (49)
Now we consider the function of $p_{j_{s}}$:
$f(p_{j_{s}})$ $=(1+\frac{|C|}{\alpha})(p_{1}^{\alpha}(\prod_{j\in C}\alpha p_{j}))^{\frac{1}{|C|+\alpha}}+p_{j_{s}}$
$\quad-(1+\frac{|C|+1}{\alpha})(p_{1}^{\alpha}(\alpha p_{j_{s}})(\prod_{j\in C}\alpha p_{j}))^{\frac{1}{|C|+1+\alpha}}.$ (50)
Then, the first derivative of $f(p_{j_{s}})$ with respect to $p_{j_{s}}$ is given by:
$f^{\prime}(p_{j_{s}})=1-\frac{1}{\alpha}(\alpha p_{1}^{\alpha}(\prod_{j\in C}\alpha p_{j}))^{\frac{1}{|C|+1+\alpha}}p_{j_{s}}^{\frac{1}{|C|+1+\alpha}-1}.$ (51)
Obviously, $f^{\prime}(p_{j_{s}})$ increases gradually as $p_{j_{s}}$ increases, and $f^{\prime}(\frac{1}{\alpha}(p_{1}^{\alpha}(\prod_{j\in C}\alpha p_{j}))^{\frac{1}{|C|+\alpha}})=0$. Since $(p_{1}^{\alpha}(\prod_{j\in C}\alpha p_{j}))^{\frac{1}{|C|+\alpha}}>\alpha p_{j_{s}}$, we can get
$f(p_{j_{s}})>f(\frac{1}{\alpha}(p_{1}^{\alpha}(\prod_{j\in C}\alpha p_{j}))^{\frac{1}{|C|+\alpha}})=0.$ (52)
From equation (52), we can observe that the value of the objective function is smaller if $j_{s}\not\in A$. In the second case, if there exists $j_{s}$ such that $(p_{1}^{\alpha}(\prod_{j\in C}\alpha p_{j}))^{\frac{1}{|C|+\alpha}}=\alpha p_{j_{s}}$ and $j_{s}\in A$, we have
$q_{j_{s}}$ $=\frac{q_{1}}{\alpha}$
$=\frac{1}{\alpha}\exp{(-1-\lambda)}(p_{1}^{\alpha}(\alpha p_{j_{s}})(\prod_{j\in C}\alpha p_{j}))^{\frac{1}{|C|+1+\alpha}}$
$=p_{j_{s}}\exp(-1-\lambda),$ (53)
which leads to the conclusion that $\mu_{j_{s}}$ = 0 according to the equation (31), contradicting the definition of the set $A$. Hence, no such $j_{s}$ can exist. In conclusion, for all $s\in\{t,t+1,\ldots,|B|\}$, it holds that $j_{s}\not\in A$.
At this time, we have successfully identified the active set $A$. Subsequently, we can obtain the optimal solution by applying equations (39), (40), and (42), which leads to the formulation of Algorithm 2.
7.2 Time Efficiency
Although the proof may be somewhat intricate, it is worth emphasizing that Algorithm 2 exhibits significantly higher efficiency in comparison to conventional convex programming tools such as MOSEK. The table below presents the average time taken by the classical convex programming tool (MOSEK) and the proposed Algorithm 2 over 1000 runs. It is evident that for $N=100$, Algorithm 2 consumes only around one-tenth of the time required by traditional convex programming tools.
Table 6: Comparisons about average training time required by Algorithm 2 and the MOSEK tool.
MOSEK Algo. 2 (Ours) Speedup ratio
Time (Avg.) 21.93 ms 2.16 ms 10.15
7.3 Further Discussion
By examining Algorithm 2, we can find that if $\alpha$ is sufficiently large such that $|A|=N-1$, the optimal soft labels degenerate into smoothed labels with a label smoothing factor of $s=N/(\alpha-1+N)$. Furthermore, as $\alpha$ increases towards infinity, the optimal soft labels converge towards one-hot labels. Then, the PCE loss degenerate into the traditional cross entropy.
8 Full Results with ResNet-50
In this section, we demonstrate the full results of our experiments that were carried out on five well-known benchmark datasets, including VLCS, PACS, OfficeHome, TerraIncognita and DomainNet. These experiments were conducted using ResNet-50, which was pre-trained on ImageNet.
8.1 VLCS
Table 7: Out-of-domain accuracies (%) on each domain of VLCS and their average.
       Algorithms        C        L        S        V        Avg.
       ERM        97.7 $\pm$ 0.4        64.3 $\pm$ 0.9        73.4 $\pm$ 0.5        74.6 $\pm$ 1.3        77.5
       IRM        98.6 $\pm$ 0.1        64.9 $\pm$ 0.9        73.4 $\pm$ 0.6        77.3 $\pm$ 0.9        78.5
       GroupDRO        97.3 $\pm$ 0.3        63.4 $\pm$ 0.9        69.5 $\pm$ 0.8        76.7 $\pm$ 0.7        76.7
       Mixup        98.3 $\pm$ 0.6        64.8 $\pm$ 1.0        72.1 $\pm$ 0.5        74.3 $\pm$ 0.8        77.4
       MLDG        97.4 $\pm$ 0.2        65.2 $\pm$ 0.7        71.0 $\pm$ 1.4        75.3 $\pm$ 1.0        77.2
       CORAL        98.3 $\pm$ 0.1        66.1 $\pm$ 1.2        73.4 $\pm$ 0.3        77.5 $\pm$ 1.2        78.8
       MMD        97.7 $\pm$ 0.1        64.0 $\pm$ 1.1        72.8 $\pm$ 0.2        75.3 $\pm$ 3.3        77.5
       DANN        99.0 $\pm$ 0.3        65.1 $\pm$ 1.4        73.1 $\pm$ 0.3        77.2 $\pm$ 0.6        78.6
       CDANN        97.1 $\pm$ 0.3        65.1 $\pm$ 1.2        70.7 $\pm$ 0.8        77.1 $\pm$ 1.5        77.5
       MTL        97.8 $\pm$ 0.4        64.3 $\pm$ 0.3        71.5 $\pm$ 0.7        75.3 $\pm$ 1.7        77.2
       SagNet        97.9 $\pm$ 0.4        64.5 $\pm$ 0.5        71.4 $\pm$ 1.3        77.5 $\pm$ 0.5        77.8
       ARM        98.7 $\pm$ 0.2        63.6 $\pm$ 0.7        71.3 $\pm$ 1.2        76.7 $\pm$ 0.6        77.6
       VREx        98.4 $\pm$ 0.3        64.4 $\pm$ 1.4        74.1 $\pm$ 0.4        76.2 $\pm$ 1.3        78.3
       RSC        97.9 $\pm$ 0.1        62.5 $\pm$ 0.7        72.3 $\pm$ 1.2        75.6 $\pm$ 0.8        77.1
       SFT (Ours)        99.5 $\pm$ 0.1        66.2 $\pm$ 0.2        74.8 $\pm$ 0.5        78.7 $\pm$ 0.4        79.8
8.2 PACS
Table 8: Out-of-domain accuracies (%) on each domain of PACS and their average.
       Algorithms        A        C        P        S        Avg.
       ERM        84.7 $\pm$ 0.4        80.8 $\pm$ 0.6        97.2 $\pm$ 0.3        79.3 $\pm$ 1.0        85.5
       IRM        84.8 $\pm$ 1.3        76.4 $\pm$ 1.1        96.7 $\pm$ 0.6        76.1 $\pm$ 1.0        83.5
       GroupDRO        83.5 $\pm$ 0.9        79.1 $\pm$ 0.6        96.7 $\pm$ 0.3        78.3 $\pm$ 2.0        84.4
       Mixup        86.1 $\pm$ 0.5        78.9 $\pm$ 0.8        97.6 $\pm$ 0.1        75.8 $\pm$ 1.8        84.6
       MLDG        85.5 $\pm$ 1.4        80.1 $\pm$ 1.7        97.4 $\pm$ 0.3        76.6 $\pm$ 1.1        84.9
       CORAL        88.3 $\pm$ 0.2        80.0 $\pm$ 0.5        97.5 $\pm$ 0.3        78.8 $\pm$ 1.3        86.2
       MMD        86.1 $\pm$ 1.4        79.4 $\pm$ 0.9        96.6 $\pm$ 0.2        76.5 $\pm$ 0.5        84.6
       DANN        86.4 $\pm$ 0.8        77.4 $\pm$ 0.8        97.3 $\pm$ 0.4        73.5 $\pm$ 2.3        83.6
       CDANN        84.6 $\pm$ 1.8        75.5 $\pm$ 0.9        96.8 $\pm$ 0.3        73.5 $\pm$ 0.6        82.6
       MTL        87.5 $\pm$ 0.8        77.1 $\pm$ 0.5        96.4 $\pm$ 0.8        77.3 $\pm$ 1.8        84.6
       SagNet        87.4 $\pm$ 1.0        80.7 $\pm$ 0.6        97.1 $\pm$ 0.1        80.0 $\pm$ 0.4        86.3
       ARM        86.8 $\pm$ 0.6        76.8 $\pm$ 0.5        97.4 $\pm$ 0.3        79.3 $\pm$ 1.2        85.1
       VREx        86.0 $\pm$ 1.6        79.1 $\pm$ 0.6        96.9 $\pm$ 0.5        77.7 $\pm$ 1.7        84.9
       RSC        85.4 $\pm$ 0.8        79.7 $\pm$ 1.8        97.6 $\pm$ 0.3        78.2 $\pm$ 1.2        85.2
       SFT (Ours)        90.1 $\pm$ 0.3        80.3 $\pm$ 1.0        98.6 $\pm$ 0.2        84.3 $\pm$ 1.4        88.3
8.3 OfficeHome
Table 9: Out-of-domain accuracies (%) on each domain of OfficeHome and their average.
       Algorithms        A        C        P        R        Avg.
       ERM        61.3 $\pm$ 0.7        52.4 $\pm$ 0.3        75.8 $\pm$ 0.1        76.6 $\pm$ 0.3        66.5
       IRM        58.9 $\pm$ 2.3        52.2 $\pm$ 1.6        72.1 $\pm$ 2.9        74.0 $\pm$ 2.5        64.3
       GroupDRO        60.4 $\pm$ 0.7        52.7 $\pm$ 1.0        75.0 $\pm$ 0.7        76.0 $\pm$ 0.7        66.0
       Mixup        62.4 $\pm$ 0.8        54.8 $\pm$ 0.6        76.9 $\pm$ 0.3        78.3 $\pm$ 0.2        68.1
       MLDG        61.5 $\pm$ 0.9        53.2 $\pm$ 0.6        75.0 $\pm$ 1.2        77.5 $\pm$ 0.4        66.8
       CORAL        65.3 $\pm$ 0.4        54.4 $\pm$ 0.5        76.5 $\pm$ 0.1        78.4 $\pm$ 0.5        68.7
       MMD        60.4 $\pm$ 0.2        53.3 $\pm$ 0.3        74.3 $\pm$ 0.1        77.4 $\pm$ 0.6        66.3
       DANN        59.9 $\pm$ 1.3        53.0 $\pm$ 0.3        73.6 $\pm$ 0.7        76.9 $\pm$ 0.5        65.9
       CDANN        61.5 $\pm$ 1.4        50.4 $\pm$ 2.4        74.4 $\pm$ 0.9        76.6 $\pm$ 0.8        65.8
       MTL        61.5 $\pm$ 0.7        52.4 $\pm$ 0.6        74.9 $\pm$ 0.4        76.8 $\pm$ 0.4        66.4
       SagNet        63.4 $\pm$ 0.2        54.8 $\pm$ 0.4        75.8 $\pm$ 0.4        78.3 $\pm$ 0.3        68.1
       ARM        58.9 $\pm$ 0.8        51.0 $\pm$ 0.5        74.1 $\pm$ 0.1        75.2 $\pm$ 0.3        64.8
       VREx        60.7 $\pm$ 0.9        53.0 $\pm$ 0.9        75.3 $\pm$ 0.1        76.6 $\pm$ 0.5        66.4
       RSC        60.7 $\pm$ 1.4        51.4 $\pm$ 0.3        74.8 $\pm$ 1.1        75.1 $\pm$ 1.3        65.5
       SFT (Ours)        65.8 $\pm$ 0.3        58.8 $\pm$ 0.3        78.3 $\pm$ 0.4        80.6 $\pm$ 0.2        70.9
8.4 TerraIncognita
Table 10: Out-of-domain accuracies (%) on each domain of TerraIncognita and their average.
      Algorithms       L100       L38       L43       L46       Avg.
      ERM       49.8 $\pm$ 4.4       42.1 $\pm$ 1.4       56.9 $\pm$ 1.8       35.7 $\pm$ 3.9       46.1
      IRM       54.6 $\pm$ 1.3       39.8 $\pm$ 1.9       56.2 $\pm$ 1.8       39.6 $\pm$ 0.8       47.6
      GroupDRO       41.2 $\pm$ 0.7       38.6 $\pm$ 2.1       56.7 $\pm$ 0.9       36.4 $\pm$ 2.1       43.2
      Mixup       59.6 $\pm$ 2.0       42.2 $\pm$ 1.4       55.9 $\pm$ 0.8       33.9 $\pm$ 1.4       47.9
      MLDG       54.2 $\pm$ 3.0       44.3 $\pm$ 1.1       55.6 $\pm$ 0.3       36.9 $\pm$ 2.2       47.7
      CORAL       51.6 $\pm$ 2.4       42.2 $\pm$ 1.0       57.0 $\pm$ 1.0       39.8 $\pm$ 2.9       47.6
      MMD       41.9 $\pm$ 3.0       34.8 $\pm$ 1.0       57.0 $\pm$ 1.9       35.2 $\pm$ 1.8       42.2
      DANN       51.1 $\pm$ 3.5       40.6 $\pm$ 0.6       57.4 $\pm$ 0.5       37.7 $\pm$ 1.8       46.7
      CDANN       47.0 $\pm$ 1.9       41.3 $\pm$ 4.8       54.9 $\pm$ 1.7       39.8 $\pm$ 2.3       45.8
      MTL       49.3 $\pm$ 1.2       39.6 $\pm$ 6.3       55.6 $\pm$ 1.1       37.8 $\pm$ 0.8       45.6
      SagNet       53.0 $\pm$ 2.9       43.0 $\pm$ 2.5       57.9 $\pm$ 0.6       40.4 $\pm$ 1.3       48.6
      ARM       49.3 $\pm$ 0.7       38.3 $\pm$ 2.4       55.8 $\pm$ 0.8       38.7 $\pm$ 1.3       45.5
      VREx       48.2 $\pm$ 4.3       41.7 $\pm$ 1.3       56.8 $\pm$ 0.8       38.7 $\pm$ 3.1       46.4
      RSC       50.2 $\pm$ 2.2       39.2 $\pm$ 1.4       56.3 $\pm$ 1.4       40.8 $\pm$ 0.6       46.6
      SFT (Ours)       57.5 $\pm$ 0.4       44.6 $\pm$ 1.4       59.6 $\pm$ 0.5       41.0 $\pm$ 1.0       50.7
8.5 DomainNet
Table 11: Out-of-domain accuracies (%) on each domain of TerraIncognita and their average.
Algorithm clip info paint quick real sketch Avg.
ERM 58.1 $\pm$ 0.3 18.8 $\pm$ 0.3 46.7 $\pm$ 0.3 12.2 $\pm$ 0.4 59.6 $\pm$ 0.1 49.8 $\pm$ 0.4 40.9
IRM 48.5 $\pm$ 2.8 15.0 $\pm$ 1.5 38.3 $\pm$ 4.3 10.9 $\pm$ 0.5 48.2 $\pm$ 5.2 42.3 $\pm$ 3.1 33.9
GroupDRO 47.2 $\pm$ 0.5 17.5 $\pm$ 0.4 33.8 $\pm$ 0.5 9.3 $\pm$ 0.3 51.6 $\pm$ 0.4 40.1 $\pm$ 0.6 33.3
Mixup 55.7 $\pm$ 0.3 18.5 $\pm$ 0.5 44.3 $\pm$ 0.5 12.5 $\pm$ 0.4 55.8 $\pm$ 0.3 48.2 $\pm$ 0.5 39.2
MLDG 59.1 $\pm$ 0.2 19.1 $\pm$ 0.3 45.8 $\pm$ 0.7 13.4 $\pm$ 0.3 59.6 $\pm$ 0.2 50.2 $\pm$ 0.4 41.2
CORAL 59.2 $\pm$ 0.1 19.7 $\pm$ 0.2 46.6 $\pm$ 0.3 13.4 $\pm$ 0.4 59.8 $\pm$ 0.2 50.1 $\pm$ 0.6 41.5
MMD 32.1 $\pm$ 13.3 11.0 $\pm$ 4.6 26.8 $\pm$ 11.3 8.7 $\pm$ 2.1 32.7 $\pm$ 13.8 28.9 $\pm$ 11.9 23.4
DANN 53.1 $\pm$ 0.2 18.3 $\pm$ 0.1 44.2 $\pm$ 0.7 11.8 $\pm$ 0.1 55.5 $\pm$ 0.4 46.8 $\pm$ 0.6 38.3
CDANN 54.6 $\pm$ 0.4 17.3 $\pm$ 0.1 43.7 $\pm$ 0.9 12.1 $\pm$ 0.7 56.2 $\pm$ 0.4 45.9 $\pm$ 0.5 38.3
MTL 57.9 $\pm$ 0.5 18.5 $\pm$ 0.4 46.0 $\pm$ 0.1 12.5 $\pm$ 0.1 59.5 $\pm$ 0.3 49.2 $\pm$ 0.1 40.6
SagNet 57.7 $\pm$ 0.3 19.0 $\pm$ 0.2 45.3 $\pm$ 0.3 12.7 $\pm$ 0.5 58.1 $\pm$ 0.5 48.8 $\pm$ 0.2 40.3
ARM 49.7 $\pm$ 0.3 16.3 $\pm$ 0.5 40.9 $\pm$ 1.1 9.4 $\pm$ 0.1 53.4 $\pm$ 0.4 43.5 $\pm$ 0.4 35.5
VREx 47.3 $\pm$ 3.5 16.0 $\pm$ 1.5 35.8 $\pm$ 4.6 10.9 $\pm$ 0.3 49.6 $\pm$ 4.9 42.0 $\pm$ 3.0 33.6
RSC 55.0 $\pm$ 1.2 18.3 $\pm$ 0.5 44.4 $\pm$ 0.6 12.2 $\pm$ 0.2 55.7 $\pm$ 0.7 47.8 $\pm$ 0.9 38.9
SFT (Ours) 64.9 $\pm$ 0.0 22.0 $\pm$ 0.3 52.5 $\pm$ 0.1 16.3 $\pm$ 0.3 64.4 $\pm$ 0.1 55.6 $\pm$ 0.4 46.0
9 Full Results with ViT-B/16 and ViT-L/14
In this section, we show the full results of visual prompt tuning with the pre-trained large-scale vision transformers (including ViT-B/16 and ViT-L/14).
9.1 VLCS
Table 12: Out-of-domain accuracies (%) on each domain of VLCS and their average.
      Backbone       Algorithms       C       L       S       V       Avg.
      ViT-B/16       ERM       95.9       66.2       81.8       79.7       80.9
      IRM       96.5       67.7       85.0       78.7       81.9
      DANN       96.6       68.3       82.1       79.8       81.7
      CDANN       95.2       66.7       85.2       80.3       81.9
      CORAL       96.6       67.4       84.3       81.5       82.5
      MMD       95.4       67.3       83.2       81.6       81.9
      IIB       97.6       65.7       84.0       82.7       82.5
      SAM       96.6       68.1       84.9       84.5       83.5
      SFT (Ours)       96.8       68.8       84.8       85.8       84.1
      ViT-L/14       ERM       95.8       66.8       86.7       82.4       82.9
      SAM       96.8       68.6       85.5       86.0       84.2
      SFT (Ours)       96.6       68.7       85.0       87.2       84.4
9.2 PACS
Table 13: Out-of-domain accuracies (%) on each domain of PACS and their average.
      Backbone       Algorithms       A       C       P       S       Avg.
      ViT-B/16       ERM       97.7       97.5       99.6       91.4       96.6
      IRM       98.2       96.7       99.8       90.7       96.4
      DANN       96.6       98.1       99.6       87.5       95.5
      CDANN       97.4       97.8       99.9       88.7       96.0
      CORAL       96.2       97.7       99.6       88.2       95.4
      MMD       96.9       97.7       99.6       86.0       95.1
      IIB       97.5       97.4       99.8       89.2       96.0
      SAM       97.8       98.1       99.7       88.9       96.1
      SFT (Ours)       98.2       98.8       99.9       90.2       96.8
      ViT-L/14       ERM       99.2       99.5       99.9       96.6       98.8
      SAM       99.3       99.6       99.9       96.0       98.7
      SFT (Ours)       99.3       99.3       99.6       96.2       98.6
9.3 OfficeHome
Table 14: Out-of-domain accuracies (%) on each domain of OfficeHome and their average.
      Backbone       Algorithms       A       C       P       R       Avg.
      ViT-B/16       ERM       83.7       73.8       89.9       89.2       84.1
      IRM       81.4       73.2       88.9       88.9       83.1
      DANN       80.7       73.1       88.9       88.2       82.7
      CDANN       82.6       71.1       87.9       87.5       82.3
      CORAL       82.7       72.9       88.4       89.3       83.3
      MMD       83.5       73.0       89.5       88.6       83.7
      IIB       81.9       73.5       90.7       89.5       83.9
      SAM       84.2       76.6       91.0       90.8       85.7
      SFT (Ours)       86.3       77.7       91.1       90.9       86.5
      ViT-L/14       ERM       89.8       83.4       93.9       93.8       90.2
      SAM       89.8       84.7       94.6       94.0       90.8
      SFT (Ours)       90.8       85.2       94.5       94.5       91.3
9.4 TerraIncognita
Table 15: Out-of-domain accuracies (%) on each domain of TerraIncognita and their average.
      Backbone       Algorithms       L100       L38       L43       L46       Avg.
      ViT-B/16       ERM       58.5       58.2       64.1       41.1       55.5
      IRM       45.7       53.5       55.4       48.8       50.9
      DANN       52.9       52.4       56.7       45.9       52.0
      CDANN       58.6       51.9       61.5       47.6       54.9
      CORAL       51.4       45.2       60.9       50.6       52.0
      MMD       57.5       57.1       62.0       50.9       56.9
      IIB       65.3       53.6       65.6       47.5       58.0
      SAM       64.6       52.0       61.5       48.3       56.6
      SFT (Ours)       70.7       58.3       65.3       50.5       61.2
      ViT-L/14       ERM       65.1       55.1       69.6       55.4       61.3
      SAM       67.3       56.4       72.8       54.8       62.8
      SFT (Ours)       65.4       61.1       71.2       62.9       65.2
9.5 DomainNet
Table 16: Out-of-domain accuracies (%) on each domain of TerraIncognita and their average.
Backbone Algorithms clip info paint quick real sketch Avg.
ViT-B/16 ERM 77.6 44.4 66.4 18.8 81.2 66.7 59.2
IRM 73.1 45.6 67.1 19.3 81.2 68.6 59.1
DANN 74.9 42.9 67.9 19.1 79.2 67.3 58.6
CDANN 74.7 44.5 66.1 19.2 79.2 67.0 58.4
CORAL 77.6 44.7 66.6 19.1 81.0 68.1 59.5
MMD 76.8 45.9 67.4 20.1 80.9 68.3 59.9
IIB 76.5 42.4 66.5 18.5 79.9 67.6 58.6
SAM 76.8 45.4 68.4 18.8 81.1 68.3 59.8
SFT (Ours) 77.9 46.3 68.6 19.4 81.6 68.9 60.5
ViT-L/14 ERM 82.7 54.5 73.1 23.7 84.3 74.4 65.4
SAM_VPT 83.1 51.9 73.3 23.1 85.3 74.7 65.2
SFT (Ours) 83.2 55.0 75.2 24.6 86.1 75.3 66.5
10 Reproducibility
To guarantee reproducibility, we will provide an explanation of the code and hyperparameters in this section.
10.1 Code
Our work is built upon DomainBed, which is released under the MIT license. All experiments are conducted on a single NVIDIA Tesla V100 or A40. Later, we will upload our code, along with the log files containing additional experimental details, to GitHub.
10.2 Hyperparameters
10.2.1 Experiments on Toy Dataset
In our toy experiments, we generate a dataset with three classes ($C_{1}$, $C_{2}$ and $C_{3}$) and four domains ($D_{1}$, $D_{2}$, $D_{3}$ and $D_{4}$). There are $3\times 100$ samples in each domain. We use data from the first three domains ($D_{1}$, $D_{2}$ and $D_{3}$) as the training domains, with the remaining domain $D_{4}$ serves as the test domain. For simplicity, we consider the covariance matrices to be diagonal with the same elements: $\boldsymbol{\Sigma}_{i}=\sigma_{i}^{2}\mathbf{I}$ and $\boldsymbol{\Sigma}_{ij}=\sigma_{ij}^{2}\mathbf{I}$. The specific parameters for each domain and class are provided in Table 17. During training, we use a linear classifier with the Adam optimizer. The batch size is set to 16 and the learning rate is 5e-4.
Table 17: Parameters for the generation of the toy dataset.
Classes $\mu_{i}$ $\sigma_{i}$ Domains $\mu_{ij}$ $\sigma_{ij}$
$C_{1}$ $(0,\sqrt{3}/2)$ 0.4 $D_{1}$ (0.71,1.03) 0.2
$D_{2}$ (-0.04,0.20) 0.2
$D_{3}$ (0.08,1.22) 0.2
$D_{4}$ (-0.52,0.54) 0.2
$C_{2}$ $(-1/2,0)$ 0.4 $D_{1}$ (-0.11,0.90) 0.2
$D_{2}$ (-0.45,0.15) 0.2
$D_{3}$ (-0.68,0.03) 0.2
$D_{4}$ (-0.81,-0.11) 0.2
$C_{3}$ $(1/2,0)$ 0.4 $D_{1}$ (1.25,-0.39) 0.2
$D_{2}$ (-0.20,0.52) 0.2
$D_{3}$ (0.80,0.23) 0.2
$D_{4}$ (0.83,-0.12) 0.2
10.2.2 Experiments on Real Dataset
The hyperparameter search spaces for ResNet-50, ViT-B/16 and ViT-L/14 are shown below. In the table, $U$ and list indicate Uniform distribution and random choice, respectively. The optimal hyperparameters and model parameters for each dataset will be available on our GitHub page.
Table 18: Hyperparameter search space for ResNet-50, ViT-B/16 and ViT-L/14.
Parameter ResNet50 ViT-B/16 ViT-L/14
batch size $U[24,32]$ $U[16,28]$ $U[8,16]$
learning rate $U[5.0e-6,5.5e-5]$ $10^{U[-4.5,-3.0]}$ $10^{U[-4.5,-3.0]}$
ResNet dropout [0.0,0.1,0.5] -- --
weight decay [1e-4, 1e-6] 0.0 0.0
$\rho$ [0.01, 0.02, 0.03, 0.05, 0.1] [0.1, 0.2, 0.3, 0.5] [0.05,0.1,0.2,0.3,0.5]
$\lambda_{1}$ $U[0,1]$ $U[0,0.5]$ $U[0,0.5]$
$\lambda_{2}$ $U[0,1]$ $U[0,0.5]$ $U[0,0.5]$
$\alpha$ $10^{U[0.5,3]}$ $10^{U[0.5,3]}$ $10^{U[0.5,3]}$
11 Broader Impacts
This paper primarily focuses on developing an effective domain generalization method to address the problem of domain shifts. Given that domain shifts are ubiquitous in real-world applications, this work has the potential to make a positive impact by learning models that are less biased towards ethical aspects. We do not foresee any significant negative social impact of this work.

### v2

@@
 On average across five DG benchmarks, SFT surpasses the sharpness-aware minimization by 2.6% with ResNet-50 and 1.5% with ViT-B/16, respectively.
-The code will be available soon.
 1 Introduction
@@
 [28] provided a detailed comparison of these methods, showing that SAM tends to lead models to wider basins, while SWA merely helps locate the center of one certain basin.
-In this paper, we focus on addressing the inconsistency of loss landscapes when SAM is applied in domain generalization (DG) scenarios.
+Several efficient variants of SAM [12, 26, 24] have also been developed for real-world applications.
+In this paper, we focus on addressing the inconsistency of loss landscapes when SAM is applied in DG scenarios.
 Loss landscapes for domain generalization.
@@
 [48] considered the domain discrepancy via formalizing a notion of consistency for minima of loss surfaces, and then proposed an algorithm (ANDMask) to alter the optimization trajectory through masking inconsistent gradient components among training domains.
-Different from all of the above methods that consider loss landscapes to be static, we attempt to refine the loss surfaces dynamically for consistency and finally seek a consistent flat minima across various domains.
+Different from all of the above methods that consider loss landscapes to be static, we attempt to refine the loss surfaces dynamically for consistency and finally seek consistent flat minima across various domains.
 3 Methodology
@@
 7      Randomly select two training domains $D_{d}$ and $D_{d^{\prime}}$.
-8      Update the model parameter $\boldsymbol{\theta}$ on the domain $D_{d}$:
-9      $\boldsymbol{\theta}\leftarrow\boldsymbol{\theta}-\eta\nabla_{\boldsymbol{\theta}}\mathcal{L}_{D_{d}}^{\mathrm{SAM-SL}}(\boldsymbol{\theta},\boldsymbol{\phi})$.
+8      Update the model parameter $\boldsymbol{\theta}$ on the domain $D_{d}$: $\boldsymbol{\theta}\leftarrow\boldsymbol{\theta}-\eta\nabla_{\boldsymbol{\theta}}\mathcal{L}_{D_{d}}^{\mathrm{SAM-SL}}(\boldsymbol{\theta},\boldsymbol{\phi})$.
 9      Calculate the feedback signal $|\mathcal{L}_{D_{d}}^{\mathrm{CES-SL}}(\boldsymbol{\theta},\boldsymbol{\phi})-\mathcal{L}_{D_{d^{\prime}}}^{\mathrm{CES-SL}}(\boldsymbol{\theta},\boldsymbol{\phi})|$ using both domains.
@@
 This paper focuses on a more fundamental problem: minimizing the “distance” between the refiner’s output and the true label space for each category.
-For instance, if the first category is the true label for the sample $\boldsymbol{x}_{i}$, the corresponding label space is defined as:
-$C_{1}=\{(q_{1},\ldots,q_{N})\mid\forall k\neq 1:q_{1}\geq\alpha q_{k},\sum_{k=1}^{N}q_{k}=1\}.$ (10)
+For instance, if the first category is the true label for the sample $\boldsymbol{x}_{i}$ (i.e., $y_{i}=1$), the corresponding label space is defined as:
+$C_{1}=\{(q_{1},\ldots,q_{N})\mid\forall k\neq 1:q_{1}\geq\alpha q_{k},\sum_{k=1}^{N}q_{k}=1\},$ (10)
 where $\alpha\geq 1$ is a hyperparameter that represents the minimum ratio between $q_{1}$ and $q_{k}$.
 Similar to the concept of “distance” in Euclidean space, we define the “distance” between the output $\tilde{\mathbf{y}}=\boldsymbol{g}_{\boldsymbol{\phi}}(\boldsymbol{x})$ and the label space $C_{1}$ as:
-$\min_{\mathbf{y}}\,\mathrm{KL}(\mathbf{y}\,||\,\tilde{\mathbf{y}})\quad\text{subject to}\quad\mathbf{y}\in C_{1}.$ (11)
+$\min_{\mathbf{y}}\,\mathrm{KL}(\mathbf{y}\,||\,\tilde{\mathbf{y}})\quad\text{subject to}\quad\mathbf{y}\in C_{1},$ (11)
+where $\mathrm{KL}(\cdot\|\cdot)$ denotes the Kullback-Leibler (KL) divergence.
 We propose an efficient algorithm (Algorithm 2) to find the optimal solution $\mathbf{y}^{\star}$ for this problem, which converges in $N$ steps ($N$ is the number of classes) and is computationally more efficient than using common convex programming tools.
@@
 Algorithm 2 KL Divergence Minimization
-To encourage the loss landscapes to become flatter while pursuing consistency, we add the the sharpness term to the previously obtained feedback signal (the sharpness difference) as an extra penalty.
+To encourage the loss landscapes to become flatter while pursuing consistency, we add the sharpness term to the previously obtained feedback signal (the sharpness difference) as an extra penalty.
 Then, the final loss function is defined as:
@@
 SFT (Ours) 79.8$\pm$0.1 88.3$\pm$0.3 70.9$\pm$0.1 50.7$\pm$0.4 46.0$\pm$0.0 67.1
-Comparison with weight average methods.
-Weight averaging methods, such as SWAD [70], seek flat minima in a different manner.
-Previous work [72] has compared it with SAM and revealed that parameters sought by the two methods are located in different basins.
-Since model averaging is a totally different method of model selection, it wouldn’t be fair to make a direct comparison with SFT.
-Nonetheless, in our experiment with ViT-B/16 (as shown in Table 4), we observed that SAM and SFT outperform SWAD by 0.6% (85.1$\rightarrow$85.7%) and 1.4% (85.1$\rightarrow$86.5%), respectively.
+4.3 Sharpness Analysis
+We also perform a quantitative analysis of the landscape consistency.
+As mentioned earlier, SFT quantifies landscape consistency through domain-specific sharpness differences.
+In Figure 4, we find that SFT reduces domain-specific sharpness values and their differences compared to SAM, while simultaneously attaining higher classification accuracy.
+These empirical results confirm that SFT effectively improves the loss landscape consistency across domains, while also demonstrating the positive impact of landscape consistency on the DG performance.
 4.4 Ablation Study
@@
 Finally, the comparison between the upper and lower sections demonstrates that the consistency-guided soft labels in our SFT framework offer a clear advantage over traditional one-hot or smoothed labels, significantly improving model generalization.
-Table 4: Comparisons of SFT and the weight average method (SWAD). We report out-of-domain accuracies with/without SWAD on four domains of OfficeHome.
-Algorithms Art Clipart Product Real World Avg.
-ERM [45] 83.7 73.8 89.9 89.2 84.1
-ERM [45]$+$SWAD [70] 84.9 74.9 90.8 89.9 85.1
-SAM [34] 84.2 76.6 91.0 90.8 85.7
-SAM [34]$+$SWAD [70] 84.6 76.5 91.2 90.8 85.8
-SFT (Ours) 86.3 77.7 91.1 90.9 86.5
-SFT (Ours) $+$ SWAD [70] 86.0 77.5 91.2 91.3 86.5
+Figure 4: Comparison of model sharpness and DG accuracy across different training strategies on the OfficeHome dataset.
 Table 4: Ablative experiments on the OfficeHome dataset.
@@
       IIB       81.9       73.5       90.7       89.5       83.9
+      SWAD       84.9       74.9       90.8       89.9       85.1
       SAM       84.2       76.6       91.0       90.8       85.7
@@
 10.1 Code
-Our work is built upon DomainBed, which is released under the MIT license.
-All experiments are conducted on a single NVIDIA Tesla V100 or A40.
-Later, we will upload our code, along with the log files containing additional experimental details, to GitHub.
+Our work is built upon DomainBed, which is released under the MIT license. All experiments are conducted on a single NVIDIA Tesla V100 or A40.
 10.2 Hyperparameters
@@
 10.2.2 Experiments on Real Dataset
-The hyperparameter search spaces for ResNet-50, ViT-B/16 and ViT-L/14 are shown below.
-In the table, $U$ and list indicate Uniform distribution and random choice, respectively.
-The optimal hyperparameters and model parameters for each dataset will be available on our GitHub page.
+The hyperparameter search spaces for ResNet-50, ViT-B/16 and ViT-L/14 are shown below. In the table, $U$ and list indicate Uniform distribution and random choice, respectively.
 Table 17: Hyperparameter search space for ResNet-50, ViT-B/16 and ViT-L/14.